### v1

# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

Abstract
[…]
1 Introduction
Figure 1: Identifiability of linear properties. Plots in the left and right dotted squares illustrate the embeddings of two next-token predictors $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$ that generate the same distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}$ within a set of conditionals distributions $p(y\mid{\bm{\mathrm{x}}})$. Theorem 5 proves a one-to-one correspondence (the dashed orange arrow) between conditional distributions and $\sim_{EL}$-equivalent models (the orange partitions of $\Theta$). This extends a result by Roeder et al. [3] characterizing $\sim_{L}$-equivalent models (green partitions of $\Theta$). Here $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ while the embedding representations are not equal up to a linear transformation (thus $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}){\color[rgb]{0.0,0.5,0.0}\definecolor[named]{pgfstrokecolor}{rgb}{0.0,0.5,0.0}\not\sim_{L}}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$), as shown by how the purple and blue parallelograms in the embeddings of the left model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ get distorted in those of the right model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. Both models display relational linear probing for the query ${\bm{\mathrm{q}}}$$=$Is the text written in English?”: one can linearly separate the embeddings of textual inputs which, when concatenated with ${\bm{\mathrm{q}}}$, have “yes” as the likeliest next token, from those which have “no”. In Proposition 14, we provide conditions under which all or none of the models in the ${\color[rgb]{1,.5,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,.5,0}\sim_{EL}}$ equivalence class share the same linear property.
In natural language processing, it is well-established that linear relationships between high-dimensional, real-valued vector representations of textual inputs reflect semantic and syntactic patterns. This was motivated in seminal works [4, 5, 6, 7, 8] and extensively validated in word embedding models [9, 10, 11] as well as modern large language models trained for next-token prediction [12, 13, 2, 14, 15, 16, 17, 18, 19].
This ubiquity is puzzling, as different internal representations can produce identical next-token distributions, resulting in distribution-equivalent but internally distinct models. This raises a key question: Are the observed linear properties shared across all models with the same next-token distribution?
Our main result is a mathematical proof that, under suitable conditions, certain linear properties hold for either all or none of the equivalent models generating a given next-token distribution. We demonstrate this through three main contributions.
The first main contribution (Section 3) is an identifiability result characterizing distribution-equivalent next-token predictors. Our result is a generalization of the main theorems by Roeder et al. [3] and Khemakhem et al. [20], relaxing the assumptions of diversity and equal representation dimensionality. This result is of independent interest for research on identifiable representation learning since our analysis is applicable to several discriminative models beyond next-token prediction [3].
Our second main contribution (Section 4) is to subsume several linear properties in a common framework. We start by defining an analogue to relational linearity [1, 2], where the definition only relies on terms appearing in our identifiability result. The key idea is to represent entities as vectors, binary relations as matrices, and to model the operation of applying a relation to an entity through matrix-vector multiplication, which yields the vector corresponding to the related entity. For example, in the sentence “Jimi Hendrix plays the guitar”, the relation between the entities “Jimi Hendrix” and “the guitar” is signified by the word “plays” and represented as a matrix-vector multiplication in representation space.111“the piano” is a less probable continuation, we only refer to the likeliest one for ease of exposition. We then define relational counterparts to linearity properties described and analyzed in previous works [21, 22, 23, 15, 19], thus making them amenable to our analysis.
Our third main contribution (Section 5) is to show that under suitable conditions, these linear properties either hold in all or none of the models generating a given distribution. For this, we combine the definitions in Section 4 and our characterization of distribution-equivalent next-token prediction models in Section 3. Thus, identifiability theory enables us to explain what linearity properties are shared across language models which are equivalent next-token predictors.
Lastly, in Section 6 we discuss implications of our findings and in Section 7 we discuss connections to related works and future research directions.
2 Preliminaries
[…]
Next-token predictors. Here, we introduce the general form of next-token predictors used in our analysis. [^2] In Appendix A we show, using a derivation by [3], that decoder-only transformer models reduce to this form. We consider models to take text sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ of tokens $\mathcal{A}$ as input. A next-token predictor $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ consists of two functions: ${\bm{\mathrm{f}}}:\mathsf{Seq}(\mathcal{A})\to\mathbb{R}^{d}$ maps sequences to their representation, called embeddings, and ${\bm{\mathrm{g}}}:\mathcal{A}\to\mathbb{R}^{d}$ maps tokens to their representation, called unembeddings. Let $\Theta_{d}$ be the set of all tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ and let $\Theta:=\bigcup_{d=1}^{\infty}\Theta_{d}$ be the set of all tuples with any dimensionality $d$. A next-token predictor models the conditional distribution of the next-token $x_{t+1}$ given the context ${\bm{\mathrm{x}}}_{1:t}$ as
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t}):=\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(x_{t+1}))}{Z({\bm{\mathrm{x}}}_{1:t})}$ (1)
where $Z({\bm{\mathrm{x}}}_{1:t}):=\sum_{y\in\mathcal{A}}\exp[{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1:t})^{\top}{\bm{\mathrm{g}}}(y)]$ is the normalizing constant. Models of the form of 1 are trained to maximize the conditional log-likelihood of the data. Provided a data distribution $p_{\mathcal{D}}$ over sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, we consider the log-likelihood:
$\mathcal{L}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}):=\mathbb{E}_{{{\bm{\mathrm{x}}}\sim}p_{\mathcal{D}}}\left[\sum_{t=1}^{T({\bm{\mathrm{x}}})-1}\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(x_{t+1}|{\bm{\mathrm{x}}}_{t:1})\right]$ (2)
where $T({\bm{\mathrm{x}}})$ indicates the length of the sequence ${\bm{\mathrm{x}}}$. For a fixed dimension $d$, the learning objective can be written as
$\max_{({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta_{d}}\;\mathcal{L}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\,.$ (3)
Since we work in the non-parameteric setting, the functions ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ are unrestricted and, thus, the expressivity of the model is determined only by $d$, the dimensionality of the representation space.
Representation dimensionality and approximation capacity. In theory and for real-valued inputs, models of the form in 1 have been proven to be universal approximators i.e., they can approximate any conditional distribution $p(x_{t+1}\mid{\bm{\mathrm{x}}}_{1:t})$ to arbitrary precision, given a sufficiently large representation dimensionality $d$ [20]; similar results may apply to next-token predictors. In practice, even if a representation dimensionality of $d$ may be sufficient to represent a given distribution well, different practitioners may choose models with representations of different dimensionality. The linear identifiability results by Khemakhem et al. [20], Roeder et al. [3] cannot be applied in this setting, since they consider models with equal representation dimensionality. Our Theorem 5 alleviates this tension between theory and practice: It characterizes identifiability of next-token predictors modeling the same conditional distribution irrespective of their representation dimensionality.
3 Identifiability of next-token predictors
We introduce a novel identifiability analysis for the model in 1. In general, a statistical model $p_{\bm{\mathrm{\theta}}}({\bm{\mathrm{x}}})$ parameterized by ${\bm{\mathrm{\theta}}}$ is said to be identifiable up to an equivalence relation $\sim$ over parameters if $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\implies{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In other words, if two parameters entail the same distribution, then they must be equivalent. The precise notion of equivalence depends on the problem setting. Although it is less commonly discussed, this implication can often be shown to hold also in the other direction so that $\sim$ is effectively a characterization of distribution-equivalent models, i.e., $p_{\bm{\mathrm{\theta}}}=p_{\tilde{\bm{\mathrm{\theta}}}}\iff{\bm{\mathrm{\theta}}}\sim\tilde{\bm{\mathrm{\theta}}}$. In this section, we define an equivalence relation over tuples $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ that characterizes models that entail the same next-token prediction distribution. In other words, we want to define an equivalence relation $\sim$ over $\Theta$ such that $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\iff({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$; we then say $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ are $\sim$-equivalent. Our characterization applies to pairs of models having different dimensionalities, i.e. $d\not=\tilde{d}$, as opposed to previous works [24, 20, 3, 25].
Previous works [24, 20, 3, 25] have shown that, under an assumption on model representations, a linear invertible transformation relates the representations of two distributionally-equivalent models. Intuitively, this assumption, known as variability [24, 20, 25] or diversity condition [3], requires that at least one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ “uses” the whole space of representations. To state the condition formally, we define the linear space spanned by the image $\text{Im}({\bm{\mathrm{h}}})\subseteq\mathbb{R}^{d}$ of a function ${\bm{\mathrm{h}}}$ as $\mathrm{SIm}({\bm{\mathrm{h}}}):=\mathrm{span}\{{\bm{\mathrm{v}}}\mid{\bm{\mathrm{v}}}\in\mathrm{Im}({\bm{\mathrm{h}}})\}$. Further, for the unembeddings, we fix an arbitrary token $y_{0}\in\mathcal{A}$ to be a pivot token for the rest of the paper and define:
${\bm{\mathrm{g}}}_{0}(y):={\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0})$ (4)
for all tokens $y\in\mathcal{A}$. The diversity condition [20, 3, 25] can then be defined as:
Definition 1 (Diversity condition).
We say a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ with representation dimensionality $d$ satisfies the diversity condition, if $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$.
Intuitively, the diversity condition states that the dimension of the spaces spanned by the output of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ match the dimension of the representation space. When both the diversity condition and $d=\tilde{d}$ hold, existing identifiability results for models similar to 1 (presented in Corollary 6) guarantee equivalence of representations up to an invertible linear transformation [20, 3]. Next, we show how to relax these two conditions via a more permissive equivalence relation.
3.1 Effective complexity of the model
We generalize previous results in a setting where the diversity condition may not hold. Starting from the conditional distribution captured by the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, we have that for every $y_{0}\in\mathcal{A}$,
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ $\propto\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))$
$\propto\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))\exp(-{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0}))$
$=\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)))\,,$
featuring that the conditional distribution ${p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})}$ is fully determined by the dot product ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$. Denote by ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ the orthogonal projectors onto $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, respectively. Since both ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ and ${\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$, the dot product between the two can be evaluated as:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}))^{\top}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ (5)
$={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)\,,$ (6)
where we used the fact that ${\bm{\mathrm{P}}}_{\mathcal{F}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{F}}$ being an orthogonal projector. In general, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ may not commute [26]. We consider the subspaces:
$\mathcal{M}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}),\quad\mathcal{N}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\perp}\,,$ (7)
which will be central to our characterization. ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ underlie, respectively, the resulting projection of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ that contributes to the dot product. Only when ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, we have that $\mathcal{M}=\mathcal{N}=\mathrm{SIm}({\bm{\mathrm{f}}})\cap\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ [26]. In general, ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}\neq{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}\neq{\bm{\mathrm{g}}}_{0}$. To see this, consider the following example.
Example 1.
Let $\{{\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2},{\bm{\mathrm{e}}}_{3}\}$ be an orthogonal basis of $\mathbb{R}^{3}$. Take $\mathcal{F}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{2})$ and $\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1},{\bm{\mathrm{e}}}_{3})$. In this case, ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ commute, so it holds $\mathcal{M}=\mathcal{N}=\mathcal{F}\cap\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{1})$.
Also, we have the following properties:
Lemma 2.
For two orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ and the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ defined on spaces of 7, it holds that (i) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}}))-\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}})-\mathrm{SIm}({\bm{\mathrm{g}}}_{0})^{\perp})$; (ii) $\mathcal{M}\subseteq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$; and (iii)
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $=\big{(}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})\big{)}^{\top}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)\,.$ (8)
All proofs can be found in Appendix B. As a consequence of this lemma, we can equivalently consider the projections ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ as the components of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}$ that are retained to evaluate the dot product of 8. This also means that whatever information is contained outside $\mathcal{M}$ and $\mathcal{N}$ is not relevant for the evaluation of the dot product. The dimensionality of $\mathcal{M}$ (and $\mathcal{N}$) can be viewed as a measure of model complexity since, intuitively, the larger $\dim(\mathcal{M})$ is, the more complex the relationship between $y$ and ${\bm{\mathrm{x}}}$ can be in $p(y\mid{\bm{\mathrm{x}}})$. For this reason, we call $\dim(\mathcal{M})$ the effective complexity of the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$.
Notice that, the effective complexity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is less than or equal to the representation dimensionality $d$.
3.2 Extended linear equivalence relation
Based on the effective complexity—which may be lower than the representation dimensionality—we introduce a weaker equivalence relation between models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$. Consider the spaces ~$\tilde{\mathcal{F}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{f}}})$ and $\tilde{\mathcal{G}}:=\mathrm{SIm}(\tilde{\bm{\mathrm{g}}}_{0})$, and define ~$\tilde{\mathcal{M}}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})$ and $\tilde{\mathcal{N}}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}})^{\perp}$.
Definition 3 (Extended linear equivalence).
Two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ are extended-linearly equivalent, if both (1) $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})$ and (2) there exist two full-rank matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{d\times\tilde{d}}$ defining, respectively, invertible transformations between $\mathcal{M}$ and $\tilde{}\mathcal{M}$, and between $\mathcal{N}$ and ~$\tilde{}\mathcal{N}$, such that ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$ and
${\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})_{\phantom{0}}$ $={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})_{\phantom{0}}$ (9)
${\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}_{0}}(y)$ $={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\,,$ (10)
for all $y\in\mathcal{A},{\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$. We denote this relation by $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$.
Figure 2: Illustration of the $\sim_{EL}$ equivalence relation. (Left) The model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ represents the embeddings over the manifold $\mathrm{Im}({\bm{\mathrm{f}}})\subsetneq\mathbb{R}^{3}$, yielding $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{3}$. To ease visualization, $\mathrm{Im}({\bm{\mathrm{f}}})$ is plotted as a continuous manifold in the figure, although this is not the case for actual textual inputs. The unembeddings lie on a two-dimensional space, $(\mathrm{SIm}({\bm{\mathrm{g}}}_{0}))\cong\mathbb{R}^{2}$, drawn in blue. Consequently the projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ map onto a two-dimensional subspace, i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{G}}$. (Right) The model $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ represents both the embeddings and the unembeddings on a two-dimensional space. We have that $\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}})=\mathrm{SIm}(\tilde{{\bm{\mathrm{g}}}_{0}})=\mathbb{R}^{2}$, which implies that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{I}}}$, thus applying these projection matrices does not modify the representations (top-right and bottom-right grids). (Center) The equivalence relation $\sim_{EL}$ specifies that both ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}$ as well as ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{{\bm{\mathrm{g}}}}$ are related by a linear invertible transformation given by the matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{3\times 2}$.
The above equivalence relation generalizes the linear equivalence already known in the literature [3, 20] to that of two models which can be related to each other on subspaces of dimension $\mathrm{dim}(\mathcal{M})\leq\min\{d,\tilde{d}\}$. It shows that, after projecting the representations to suitable equal-dimensional subspaces, namely $\mathcal{M},\mathcal{N},\tilde{}\mathcal{M}$, and ~$\tilde{}\mathcal{N}$, we can find an invertible linear transformation relating them. Here, the dimension of $\mathcal{M}$ and $\tilde{}\mathcal{M}$ are equal, requiring that two equivalent models share the same effective complexity. Furthermore, the condition between ${\bm{\mathrm{M}}}$ and ${\bm{\mathrm{N}}}$ preserves the dot-product:
Proposition 4.
If $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$, then
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)=\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$ (11)
As a consequence, models in the $\sim_{EL}$ equivalence class also satisfy $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. We show that it is an equivalence relation in Section B.3 and the explicit form of ${\bm{\mathrm{N}}}$ to make the definition hold. An example of this equivalence relation is illustrated in Figure 2.
3.3 Identifiability of next-token predictors
We can relate the representations of two models based on this equivalence relation when they express the same conditional distribution. The following theorem offers a characterization of models attaining the same conditional probability distribution:
Theorem 5.
For all $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta$, which may have different representation dimensions $d$ and $\tilde{d}$, we have
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}\iff({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\,.$ (12)
With Theorem 5, we characterize the space of distributionally-equivalent next-token predictors. Models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ in the equivalence relation $\sim_{EL}$ are associated to a single conditional probability distribution $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$. This is a one-to-one correspondence between the set of conditional probability distributions expressed by 1 and the set of equivalence classes of $\sim_{EL}$ (cf. Figure 1).
Interestingly, this characterization highlights the fact that our notion of effective complexity, which we defined for models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, is a well-defined complexity measure for distributions $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}$ in the sense that it does not depend on the specific choice of parameterization. Indeed, the result implies that
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}\implies\dim(\mathcal{M})=\dim(\tilde{}\mathcal{M})\,.$
Furthermore, We can show that the previous theorem reduces to linear identifiability when both the diversity condition and $d=\tilde{d}$ hold:
Corollary 6 (Adapted from Roeder et al. [3]).
For all $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta_{d}$ such that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ satisfies the diversity condition (Definition 1), we have
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}\implies({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{L}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\,,$ (13)
where, by definition, $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{L}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ if and only if there exists an invertible matrix ${\bm{\mathrm{M}}}\in\mathbb{R}^{d\times d}$ such that for all $y\in\mathcal{A},{\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ we have
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})\ \text{and}\ {\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{M}}}^{-\top}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$
Implications for trained models.
Now assume that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta_{d}$ and $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\in\Theta_{\tilde{d}}$ are both maximizers of 3. If both models have enough capacity to represent the ground-truth data distribution $p_{\mathcal{D}}$, then they necessarily represent the same distribution, i.e. $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\mathcal{D}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}$. By Theorem 5, we can thus conclude that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{{\bm{\mathrm{g}}}})$. Assuming these assumptions hold in practice, this would imply that all models trained with sufficient capacity on a given data distribution $p_{\mathcal{D}}$ will be $\sim_{EL}$-equivalent. This analysis relies crucially on the assumption that $\Theta$ has enough capacity to represent $p_{\mathcal{D}}$. In practice, this might not hold for at least two reasons. First, we typically train with a fixed representation dimension $d$ which limits the expressivity of the model. Moreover, despite the universal approximation guarantee, there might not even exist a sufficiently large $d$ such that the model can express $p_{\mathcal{D}}$ exactly. Secondly, all distributions that can be represented by $\Theta$ put nonzero probability mass on all text sequences (because of the $\exp$), whereas one can argue that the ground-truth distribution of text $p_{\mathcal{D}}$ gives zero probability to many sequences.
4 Linear properties
We will now focus on linear properties of language models. These describe how a given model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ represents different inputs—similar to our opening example, illustrating a geometric relationship (parallelism) among the vector differences between the embeddings of two different inputs (“lucky” and “luckiest”) and that between two further inputs (“easy” and “easiest”). These linear properties are not to be confused with the linear equivalence class $\sim_{L}$ mentioned in Corollary 6, which instead describes how different models represent the same data distribution. In Section 5, we will combine the linear properties defined here with our equivalence class of next-token predictors, described in Section 3, to study what properties are shared by all models within the same equivalence class. See also Figure 1 for an illustration.
Context-query-reply sequences.
To describe linear properties, we use a set up similar to Paccanaro and Hinton [1], Hernandez et al. [2]. We consider sequences ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ that can be decomposed as ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$, where ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ denotes the context (or subject), ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ denotes the query (or relation), and $y\in\mathcal{A}$ is the reply (or object). The following example showcases the setting.
Example 2.
We consider sentences of the form ${{\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y}$ where, for example, ${\bm{\mathrm{s}}}=$“All roads lead to Rome”, ${\bm{\mathrm{q}}}=$“What is the written language?”, with most likely prediction $y=$“English”. This resembles a setup where the model is tested to determine whether it understands the property of “written language”. Another example could be ${\bm{\mathrm{s}}}=$“Rome”, ${\bm{\mathrm{q}}}=$“is the capital of” and $y=$“Italy”.
Sentences ${\bm{\mathrm{x}}}={\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$ like these can be found in natural language, capturing relational aspects between ${\bm{\mathrm{s}}}$ and $y$, mediated by ${\bm{\mathrm{q}}}$. In our setting, though, formally these tuples allow for arbitrary ${\bm{\mathrm{s}}}$ and ${\bm{\mathrm{q}}}$, including queries whose expected reply is independent of the context, like ${\bm{\mathrm{q}}}=$Whatever follows reply with $42$.”, or paraphrases, e.g., ${\bm{\mathrm{q}}}^{\prime}=$Now reply with $42$!”. In Appendix E, we demonstrate how our analsys meaningfully captures these corner cases.
4.1 Relational linear property
We describe the linear structure of representations solely relationally, that is, in terms of context-query-reply sequences ${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$, similar to Paccanaro and Hinton [1]. Relatedly, Hernandez et al. [2] provided empirical evidence for relational linearity between hidden transformer layers.
Our analysis focuses on embeddings ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})\in\mathrm{SIm}({\bm{\mathrm{f}}})$ and subspaces $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, allowing us to define relational linear properties solely in terms of the quantities described in our identifiability result; this enables our analysis to be agnostic to assumptions on the data generating process underlying natural language and, e.g., to not require positing nor referring to unobserved variables.
Definition 7 ($\Gamma$lr: Relational linearity of ${\bm{\mathrm{q}}}$ in $\Gamma$).
For a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$, let $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ be a subspace. We say that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents the query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ on $\Gamma$, if, there exist a matrix ${\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d\times d}$ and a vector ${\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\in\mathbb{R}^{d}$ such that, for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$,
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}\big{(}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\big{)}.$ (14)
When this holds, we indicate with $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$.
Intuitively, all the information relevant within ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ to compute the probability of tokens in $\Gamma$ is captured, up to an affine transformation, by ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ in the subspace $\Gamma_{\bm{\mathrm{q}}}$. Indeed, one can show that, if ${\bm{\mathrm{g}}}_{0}({\bm{\mathrm{y}}})\in\Gamma$, then it must be that ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)=\big{(}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\big{)}^{\top}{\bm{\mathrm{g}}}_{0}(y)$. The space $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$ are central to proving whether relational linearity holds for all $\sim_{EL}$-equivalent models, as we will show in Section 5.
4.2 Connection to other linear properties
Here, we show how to specialize the notion of relational linearity to capture three linear properties that have been empirically observed. We follow the taxonomy by Park et al. [23]. All proofs are reported in Appendix C.
4.2.1 Linear subspaces (ls)
[…]
Definition 8 (Parallelism in $\Gamma$).
We say that two vectors ${\bm{\mathrm{\gamma}}},{\bm{\mathrm{\gamma}}}^{\prime}\in\mathbb{R}^{d}$ are parallel in $\Gamma$ if there exists $\beta\neq 0$ (see Remark 20) such that ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{\gamma}}}=\beta\cdot{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{\gamma}}}^{\prime}$.
We show next that parallel vectors induce similar log ratios of conditional probabilities, as noted in [23, 28]:
Lemma 9.
Take a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. For $y_{0},y_{1},y_{2},y_{3}\in\mathcal{A}$, the difference vectors ${\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0})$ and ${\bm{\mathrm{g}}}(y_{3})-{\bm{\mathrm{g}}}(y_{2})$ are parallel in $\mathcal{N}$ if and only if there exists $\beta\neq 0,\text{ s.t. }\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ it holds that
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}})}=\beta\cdot\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{2}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{3}\mid{\bm{\mathrm{s}}})}\,.$ (15)
That is, parallel difference vectors for token pairs correspond to proportional likelihood ratios between the tokens in each token pair. Parallel vectors can be associated with their one-dimensional subspace, and are possibly related to changes in replies for a query ${\bm{\mathrm{q}}}$. We can capture this through the following definition:
Definition 10 (ls: Relational Linear Subspaces).
A model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ linearly represents a subspace ${\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$ relative to ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ if, for all couples of tokens $y_{i},y_{j}\in\mathcal{A}$ such that ${\bm{\mathrm{g}}}_{i}(y_{j}):={\bm{\mathrm{g}}}(y_{j})-{\bm{\mathrm{g}}}(y_{i})\in\Gamma$, there exists a vector ${\bm{\mathrm{\gamma}}}\in\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ such that $\forall{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ it holds that:
${\bm{\mathrm{g}}}_{i}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})={\bm{\mathrm{\gamma}}}^{\top}({\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})-{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}).$ (16)
An example is visualized in Figure 3. We show that the ls property is implied by relational linearity in the following sense:
[…]
Suppose that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (1) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, and also (2) ${\Gamma_{\bm{\mathrm{q}}}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})}$, then the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents $\Gamma_{\bm{\mathrm{q}}}$ relative to ${\bm{\mathrm{q}}}$ (Definition 10).
[…]
4.2.2 Linear probing (lp)
There is empirical evidence for sentence embeddings being linearly separable into embeddings of text of different languages [23, 29]. Such a structure also enables what is called linear probing [30, 31] (see below). We define linear probing as relational property:
Definition 12 (lp: Relational Linear Probing).
We say that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ can be linearly probed for a query ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ and a collection $\mathcal{Y}_{P}\subseteq\mathcal{A}$ of $\ell$ elements if there exist ${\bm{\mathrm{W}}}\in\mathbb{R}^{\ell\times d}$ and ${\bm{\mathrm{b}}}\in\mathbb{R}^{\ell}$, such that for all ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ and $\forall i\in[\ell]$ it holds that
$\mathrm{softmax}\big{(}{\bm{\mathrm{W}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{b}}}\big{)}_{i}=p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P}),$ (17)
where $p(y\mid\cdot\,;\mathcal{Y}_{P})=p(y\mid\cdot\,)/(\sum_{y^{\prime}\in\mathcal{Y}_{P}}p(y^{\prime}\mid\cdot\,))$ denotes the conditionals restricted to $\mathcal{Y}_{P}$.
To illustrate why this is termed linear probing, suppose a model given the query ${\bm{\mathrm{q}}}=$Is the text written in English?” discriminates input sequences ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ between positive $y_{0}=$yes” and negative examples $y_{1}=$no”, that is it assigns high probability mass to either $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ or to $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$. Then, these conditional distributions can be evaluated directly from ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ via a linear probe. See Figure 1 for an illustration. We show that $\Gamma$lr allows for lp:
[…]
If a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ (1) linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, and (2) ${\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j})\in\Gamma$ for all $y_{i}\in\mathcal{Y}_{P}$, then the model can be linear probed (Definition 12) for ${\bm{\mathrm{q}}}$ and $\mathcal{Y}_{P}$, with parameters given by ${\bm{\mathrm{W}}}=({\bm{\mathrm{w}}}_{1},\ldots,{\bm{\mathrm{w}}}_{\ell})^{\top}$ and ${\bm{\mathrm{b}}}=(b_{1},\ldots,b_{\ell})^{\top}$, where ${\bm{\mathrm{w}}}_{i}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and $b_{i}:=({\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})^{\top}{\bm{\mathrm{g}}}(y_{i})$.
4.2.3 Linear Steering
Similar to LP, in Section C.1 we discuss a relational counterpart to linear properties discussed by Park et al. [23], Hernandez et al. [2].
5 Linear properties shared by all distribution-equivalent models
Based on Theorem 5, we analyze which linear properties are shared across models expressing the same conditional distribution. To this end, pick a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$ (Definition 7), and also consider the space $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})$. We show that under an additional condition on $\Gamma$ and $\Gamma_{\bm{\mathrm{q}}}$, two models that are $\sim_{EL}$-equivalent share the same linear properties (all proofs are in Appendix D):
Proposition 14.
For two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\in\Theta$, such that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, it holds: if ${\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma\subseteq\mathcal{N}$, and moreover, $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$, then $\tilde{\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\tilde{\Gamma}\subseteq\tilde{}\mathcal{N}$, where $\tilde{\Gamma}=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma})$ and ${\bm{\mathrm{N}}}$ is the matrix relating ${\bm{\mathrm{g}}}_{0}$ and $\tilde{{\bm{\mathrm{g}}}}_{0}$ by the equivalence relation in Definition 3.
This shows how the relational linearity ($\Gamma$lr), combined with the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$ is a property of all or none next-token predictors modeling the same conditional distribution. As a consequence, the same holds for ls (by Proposition 11) and lp (by Proposition 13). Intuitively, the extra condition underlies that relational linearity of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ is displayed by the components of ${\bm{\mathrm{f}}}$ that contribute to the dot product with ${\bm{\mathrm{g}}}_{0}$. A $\sim_{EL}$-equivalent model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ would linearly transform these components, thus preserving relational linearity. Vice versa, since all components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$ can be arbitrarily distorted, any property of $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ that depends on those components may not hold for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$. The extra condition precisely avoids that. [^3] If $\Gamma\not\subseteq\mathcal{N}$, then relational linearity (Definition 7) would be trivially satisfied for $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and, in turn, also for $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, because ${\bm{\mathrm{\gamma}}}\in\Gamma$ would give ${\bm{\mathrm{\gamma}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=0$, and so ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\mathbf{0}$. Notice that the special case where the diversity condition (Definition 1) holds implies a similar conclusion because the condition that $\Gamma\subseteq\mathcal{N}$ and $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathcal{M}$ is then always satisfied (as $\mathcal{M}=\mathcal{N}=\mathbb{R}^{d}$).
In contrast, parallel vectors may not be preserved: Two parallel vectors in one model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ may not be parallel in another model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ with the same conditional distribution. They remain parallel only within the subspaces $\mathcal{N}$ and ~$\tilde{}\mathcal{N}$, respectively:
Proposition 15.
For two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\in\Theta$, such that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$, the vectors ${\bm{\mathrm{\gamma}}},{\bm{\mathrm{\gamma}}}^{\prime}\in\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ are parallel within $\mathcal{N}$ if and only if the corresponding vectors $\tilde{{\bm{\mathrm{\gamma}}}},\tilde{{\bm{\mathrm{\gamma}}}}^{\prime}\in\mathrm{SIm}(\tilde{\bm{\mathrm{g}}}_{0})$ are parallel in ~$\tilde{}\mathcal{N}$.
6 Discussion
Proposition 14 gives an example of a property that all distribution-equivalent next-token predictors, as characterized by our identifiability result (Theorem 5), must share. For the widely observed linear properties of language models, one may then ask whether they are indeed examples of shared properties akin to the one in Proposition 14. Tautologically, claims about the ubiquity of linear properties cannot solely be based on observations of linearity in individual model instances. One might therefore expect that only those properties shared across all equivalent models should be ubiquitously observable. [^4] An analogy could be made with the principle of covariance in physics [32, 33], which asserts that physical laws should be expressible as coordinate-independent and reference-frame-independent geometric relationships between objects that represent physical entities [34]. Recently, Villar et al. [35] suggested that this principle could inspire future developments in machine learning. In the following, we critically examine this hypothesis in light of empirical evidence.
Do we observe properties which are not shared across all $\sim_{EL}$-equivalent models?
Proposition 15 shows that vector parallelism is only preserved within a linear subspace $\mathcal{N}$ of the unembedding space. Surprisingly, a different kind of parallelism, which according to our theory is not shared by all $\sim_{EL}$-equivalent models, appears to be consistently observed in language models. In fact, several empirical studies apply dimensionality reduction through PCA to the embeddings and unembeddings to reveal and visualize linear properties including parallelism (e.g., [9, Figure 2]; [15, Figure 1]). Note that, across $\sim_{EL}$-equivalent models, the embeddings and unembeddings may not be completely contained within $\mathcal{M}$ and $\mathcal{N}$, respectively: that is the case when both $\mathcal{M}\subsetneq\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{N}\subsetneq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. As a consequence, unbounded distortions within the orthogonal complements of $\mathcal{M}$ and of $\mathcal{N}$ are inconsequential for the dot product between the embedding and the unembedding vectors (see, e.g.Figure 2, top left, for an embedding manifold not contained within any proper linear subspace). If these distortions were sufficiently large, they would prevent the visualization of vectors parallel in the sense of Proposition 15 through PCA, as the distortions would dominate the covariance matrix on which the PCA of the representations is performed, and thus the first principal components would mostly reflect those. This suggests that, in models where PCA reveals parallelism, these distortions are small, and the representations live close to a proper linear subspace.
How can we explain this? These observations suggest that something other than the assumptions of our Theorem 5 determines what models are learned in practice. One possible explanation is that some additional assumptions and constraints are at play which imply that only models in a subset of the $\sim_{EL}$ equivalence class are observed empirically. For parallelism, this could occur, for example, if the modeler chooses a fixed $d$ for which the diversity condition happens to hold (Definition 1): in which case, the resulting equivalence class would be $\sim_{L}$, and parallelism in $\mathbb{R}^{d}$ (Definition 8) is a shared property across $\sim_{L}$-equivalent models with representation dimensionality $d$. An alternative possibility is that other inductive biases, not captured by the identifiability result, are influencing the learned representations. These biases could stem from the training algorithm or architecture, steering the model toward a subset of the $\sim_{EL}$-equivalent models. Our contribution is to provide a mathematical framework that enables a clear articulation of these questions, guiding future empirical investigation.
7 Related work and future directions
Linear properties of next-token predictors have attracted widespread attention, also beyond language modeling [36, 37, 38]. More complex, non-linear properties have also been observed, such as circular token representations [39]. Formalizing these properties and investigating whether all distribution-equivalent next-token predictors share them, in the sense we studied for linear properties, is an interesting open venue.
Theoretical studies on linear properties. Park et al. [23] introduce binary latent concepts to describe several linear properties (though not relational linearity) in a unified framework. This was also applied to study categorical and hierarchical concepts [40]. Jiang et al. [28] explain linear properties of language models based on assumptions on the data-generating process and latent variables underlying natural text. This allows them to reason about the origins of linearity; in this work, we instead focus on the ubiquity of linearity in the models, taking an agnostic stance on latent concept variables. An exciting direction for future work is to prove, within our framework, why and how linear properties emerge, if they do at all.
Identifiability of representations is a central theme in generative modeling, [41, 42], particularly in non-linear ICA, e.g., [43, 44, 45, 46, 47]; and in causal representation learning, e.g., [48, 49, 50, 51, 52, 53, 54]. Reizinger et al. [55] discuss the role identifiability may play in explaining several properties of large language models [56]. Our work highlights the role identifiability plays in explaining the ubiquity of linear properties.
Acknowledgments
We thank Beatrix Miranda Nielsen, Antonio Vergari, Frederik Hytting Jørgensen, Filippo Camilloni, Adrián Javaloy, and Julius von Kügelgen for useful discussions. We acknowledge positive feedback by Stefano Teso and interesting conversations with the participants of the 2024 Bellairs Workshop on Causality. E.M. acknowledges support from TANGO, Grant Agreement No. 101120763. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Health and Digital Executive Agency (HaDEA). Neither the European Union nor the granting authority can be held responsible for them. L.G. was supported by the Danish Data Science Academy (DDSA), Grant ID: 2023-1250.
Appendices
[…]
Appendix A Functional Form of Decoders-only Transformers
Decoders-only transformer models can be reduced to the form we use in the main text, as already shown in a derivation by Roeder et al. [3]. We consider autoregressive, GPT-like models [57, 58], focusing on the GPT-J model [59].
We denote with ${\bm{\mathrm{h}}}({\bm{\mathrm{x}}};\theta)$ the representation given by the a transformer model ${\bm{\mathrm{h}}}$ with trainable parameters $\theta\in\mathbb{R}^{q}$ for the architecture. We consider the transformer to represent sentences of lenght up to $C$, and by convention we pick the last $C$ tokens if the sentence has length $t({\bm{\mathrm{x}}})>C$. Denote with $\tau:=\max(t({\bm{\mathrm{x}}})-C,1)$ it holds that:
${\bm{\mathrm{h}}}({\bm{\mathrm{x}}};\theta)={\bm{\mathrm{h}}}({\bm{\mathrm{x}}}_{t({\bm{\mathrm{x}}})-\tau:t({\bm{\mathrm{x}}})};\theta)\in\mathbb{R}^{d\times\tau}\,.$ (18)
We focus on the representation of the last token of sentence, which is used to perform next-token prediction, and denote it with
${\bm{\mathrm{h}}}({\bm{\mathrm{x}}};\theta)_{-1}\in\mathbb{R}^{d}\,.$ (19)
When predicting the next token among $K:=|\mathcal{A}|$ possible options, a common approach is to use a layer (or head) with weights ${\bm{\mathrm{W}}}_{D}\in\mathbb{R}^{K\times d}$ and a bias ${\bm{\mathrm{b}}}\in\mathbb{R}^{K}$. The representation of ${\bm{\mathrm{g}}}_{\varphi}$ is determined by the parameters $\varphi=({\bm{\mathrm{W}}}_{D},{\bm{\mathrm{b}}})$, which are used to compute the logits as follows:
$\mathrm{logits}({\bm{\mathrm{x}}}):={\bm{\mathrm{W}}}_{D}{\bm{\mathrm{h}}}({\bm{\mathrm{x}}};\theta)_{-1}+{\bm{\mathrm{b}}}\,.$ (20)
By extending the representation of ${\bm{\mathrm{h}}}_{-1}$ to ${\bm{\mathrm{f}}}_{\theta}({\bm{\mathrm{x}}}):=(1,{\bm{\mathrm{h}}}({\bm{\mathrm{x}}};\theta)^{\top}_{-1})^{\top}$, we can incorporate the bias by adding one column to ${\bm{\mathrm{W}}}_{D}$, obtaining:
$\mathrm{logits}({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{W}}}}_{D}{\bm{\mathrm{f}}}_{\theta}({\bm{\mathrm{x}}}),\quad\text{where }\tilde{{\bm{\mathrm{W}}}}_{D}=\begin{pmatrix}{\bm{\mathrm{b}}},{\bm{\mathrm{W}}}_{D}\end{pmatrix}\,.$ (21)
To predict the probability of the next token $y$ it is then sufficient to consider
$\log p_{{\bm{\mathrm{f}}}_{\theta},{\bm{\mathrm{g}}}_{\varphi}}(y\mid{\bm{\mathrm{x}}})=\mathrm{logits}({\bm{\mathrm{x}}})_{y}-\log\sum_{y^{\prime}\in\mathcal{A}}\mathrm{logits}({\bm{\mathrm{x}}}_{y^{\prime}})\,.$ (22)
Transforming the token $y\in\mathcal{A}$ to its one-hot representation, ${\bm{\mathrm{y}}}\in\{0,1\}^{K}$, we can write
${\bm{\mathrm{g}}}_{\varphi}(y)=\tilde{{\bm{\mathrm{W}}}}_{D}^{\top}{\bm{\mathrm{y}}},$ (23)
thereby leading to the expression:
$\log p_{{\bm{\mathrm{f}}}_{\theta},{\bm{\mathrm{g}}}_{\varphi}}(y\mid{\bm{\mathrm{x}}})={\bm{\mathrm{g}}}_{\varphi}(y)^{\top}{\bm{\mathrm{f}}}_{\theta}({\bm{\mathrm{x}}})-\log\sum_{y^{\prime}\in\mathcal{A}}{\bm{\mathrm{g}}}_{\varphi}(y)^{\top}{\bm{\mathrm{f}}}_{\theta}({\bm{\mathrm{x}}}),$ (24)
where we used:
${\bm{\mathrm{f}}}_{\theta}({\bm{\mathrm{x}}})=\begin{pmatrix}1\\ {\bm{\mathrm{h}}}({\bm{\mathrm{x}}}_{t({\bm{\mathrm{x}}})-\tau:t({\bm{\mathrm{x}}})};\theta)_{-1}\end{pmatrix},\qquad{\bm{\mathrm{g}}}_{\varphi}(y)=\tilde{{\bm{\mathrm{W}}}}_{D}^{\top}{\bm{\mathrm{y}}}.$ (25)
This explicit form of the embedding and unembedding respectively and consistent with (1), as previously detailed by Roeder et al. [3].
Appendix B Proofs of Section 3
B.1 Reminder of useful properties of pseudo-inverses
We will often make use of the pseudo-inverse ${\bm{\mathrm{A}}}^{+}$ of a matrix ${\bm{\mathrm{A}}}$ [60, page 221]. We denote with ${\bm{\mathrm{T}}}\mid_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}$ the restriction of ${\bm{\mathrm{T}}}$ to its orthogonal complement of the kernel [60].
Definition 16 (Pseudo-inverse).
Let ${\bm{\mathrm{T}}}\in\mathbb{R}^{m\times n}$ be a matrix. The pseudo-inverse ${\bm{\mathrm{T}}}^{+}\in\mathbb{R}^{n\times m}$ of ${\bm{\mathrm{T}}}$ is defined as the linear map:
${\bm{\mathrm{T}}}^{+}{\bm{\mathrm{w}}}:=\big{(}{\bm{\mathrm{T}}}\mid_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}\big{)}^{-1}{\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{T}}})}{\bm{\mathrm{w}}}$ (26)
for all ${\bm{\mathrm{w}}}\in\mathbb{R}^{n}$.
Accordingly, the pseudo-inverse always exists and it is unique. Notice that for any matrix ${\bm{\mathrm{T}}}\in\mathbb{R}^{m\times n}$ it holds:
${\bm{\mathrm{T}}}{\bm{\mathrm{T}}}^{+}$ $={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{T}}})}$ (27)
${\bm{\mathrm{T}}}^{+}{\bm{\mathrm{T}}}$ $={\bm{\mathrm{P}}}_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}$ (28)
${\bm{\mathrm{T}}}^{\top}({\bm{\mathrm{T}}}^{\top})^{+}$ $={\bm{\mathrm{P}}}_{\mathrm{ker}({\bm{\mathrm{T}}})^{\perp}}$ (29)
$({\bm{\mathrm{T}}}^{\top})^{+}{\bm{\mathrm{T}}}^{\top}$ $={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{T}}})}\,.$ (30)
B.2 Proof of Lemma 2
We provide here a longer version of the Lemma capturing different properties between the projectors.
Lemma (Ref Lemma 2).
Let $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$, and take $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. For the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ and the orthogonal projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ projecting on the spaces:
$\mathcal{M}=\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}),\quad\mathcal{N}=\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\perp}$ (31)
it holds:
  • (i)
    $\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathcal{F})-\mathrm{dim}\big{(}\mathcal{F}\cap\mathcal{G}^{\perp}\big{)}$;
  • (ii)
    The orthogonal projectors are also given by:
    ${\bm{\mathrm{P}}}_{\mathcal{M}}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+},$ $\quad{\bm{\mathrm{P}}}_{\mathcal{N}}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})$ (32)
    ${\bm{\mathrm{P}}}_{\mathcal{M}}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}),$ $\quad{\bm{\mathrm{P}}}_{\mathcal{N}}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}$ (33)
  • (iii)
    We have:
    ${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\phantom{+}}=$ $({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\phantom{+}}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\phantom{+}}{\bm{\mathrm{P}}}_{\mathcal{N}}$ (34)
    ${\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}=$ $({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}{\bm{\mathrm{P}}}_{\mathcal{M}}$ (35)
    ${\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{\phantom{+}}=$ $({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{\phantom{+}}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{\phantom{+}}{\bm{\mathrm{P}}}_{\mathcal{M}}$ (36)
    ${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}=$ $({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}{\bm{\mathrm{P}}}_{\mathcal{N}}$ (37)
  • (iv)
    $\mathcal{M}\subseteq\mathcal{F}$ and $\mathcal{N}\subseteq\mathcal{G}$;
  • (v)
    ${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{G}}$;
  • (vi)
    It holds ${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}$ and in particular:
    ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)=({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}))^{\top}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ (38)
Proof.
(i) By the rank-nullity theorem [60, page 62], we have that $\mathrm{dim}\,\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=d-\mathrm{dim}\,\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})$. Notice that $\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathbb{R}^{d}\setminus\mathcal{N}$, therefore
$\mathrm{dim}(\mathcal{M})=\mathrm{dim}\,\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=d-\mathrm{dim}\,\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=d-d+\mathrm{dim}\,\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\perp}=\mathrm{dim}(\mathcal{N})\,.$ (39)
Next, we derive the dimensionality of $\mathcal{M}$. Recall that for two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [60, page 62], it follows that:
$\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})=\mathrm{rank}({\bm{\mathrm{B}}})-\mathrm{dim}(\mathrm{ker}({\bm{\mathrm{A}}})\cap\mathrm{Im}({\bm{\mathrm{B}}}))\,.$ (40)
Using this for ${\bm{\mathrm{A}}}={\bm{\mathrm{P}}}_{\mathcal{G}}$ and ${\bm{\mathrm{B}}}={\bm{\mathrm{P}}}_{\mathcal{F}}$, we have that:
$\mathrm{dim}(\mathcal{M})$ $=\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})$ (41)
$=\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{F}})-\mathrm{dim}(\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{G}})\cap\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}))$ (42)
$=\mathrm{dim}(\mathcal{F})-\mathrm{dim}(\mathcal{G}^{\perp}\cap\mathcal{F})$ (43)
$=\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{f}}}))-\mathrm{dim}(\mathrm{SIm}({\bm{\mathrm{g}}}_{0})^{\perp}\cap\mathrm{SIm}({\bm{\mathrm{f}}}))\,.$ (44)
(ii) From the property of the pseudo-inverse, see 27 and 28, we have that
$({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})}={\bm{\mathrm{P}}}_{\mathcal{M}}$ (45)
and that:
$({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})={\bm{\mathrm{P}}}_{\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\perp}}={\bm{\mathrm{P}}}_{\mathcal{N}}\,.$ (46)
From 45, taking the transpose of ${\bm{\mathrm{P}}}_{\mathcal{M}}$ we get:
${\bm{\mathrm{P}}}_{\mathcal{M}}^{\top}$ $=(({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+})^{\top}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\top}$ (47)
${\bm{\mathrm{P}}}_{\mathcal{M}}$ $=({\bm{\mathrm{P}}}_{\mathcal{G}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}^{\top})^{+}({\bm{\mathrm{P}}}_{\mathcal{G}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}^{\top})$ (48)
$=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})\,,$ (49)
and from 46, taking the transpose we obtain:
${\bm{\mathrm{P}}}_{\mathcal{N}}^{\top}$ $=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\top}(({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+})^{\top}$ (50)
${\bm{\mathrm{P}}}_{\mathcal{N}}$ $=({\bm{\mathrm{P}}}_{\mathcal{G}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}^{\top})({\bm{\mathrm{P}}}_{\mathcal{G}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}^{\top})^{+}$ (51)
$=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}\,.$ (52)
(iii) Denote ${\bm{\mathrm{A}}}:={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$. From the pseudo-inverse definition [60, page 221], it holds
$({\bm{\mathrm{A}}}{\bm{\mathrm{A}}}^{+}){\bm{\mathrm{A}}}={\bm{\mathrm{A}}}={\bm{\mathrm{A}}}({\bm{\mathrm{A}}}^{+}{\bm{\mathrm{A}}})$ (53)
and substituting for ${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$ and the projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{N}}$ we get:
${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}){\bm{\mathrm{P}}}_{\mathcal{N}}\,.$ (54)
Similarly, for the pseudo-inverse, consider
$({\bm{\mathrm{A}}}^{+}{\bm{\mathrm{A}}}){\bm{\mathrm{A}}}^{+}={\bm{\mathrm{A}}}^{+}={\bm{\mathrm{A}}}^{+}({\bm{\mathrm{A}}}{\bm{\mathrm{A}}}^{+})$ (55)
and substituting we get
${\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}{\bm{\mathrm{P}}}_{\mathcal{M}}\,.$ (56)
Similarly to point (ii), taking the transpose of 54, we obtain:
$({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}){\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})\,,$ (57)
and taking the transpose of 56 we obtain:
$({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}{\bm{\mathrm{P}}}_{\mathcal{N}}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}={\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}\,.$ (58)
(iv) To show this, notice that $\mathcal{M}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})=\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{\perp}$. Therefore, we have that
$\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{\perp}=\{{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{v}}}\not\in\mathcal{G}^{\perp}\mid{\bm{\mathrm{v}}}\in\mathbb{R}^{d}\},\quad\mathcal{F}=\{{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{v}}}\neq\mathbf{0}\mid{\bm{\mathrm{v}}}\in\mathbb{R}^{d}\}\,,$ (59)
which shows that $\mathcal{M}$ is a subset of $\mathcal{F}$, and they are equal only when ${\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{I}}}$: $\mathcal{M}\subseteq\mathcal{F}$. Similarly, for $\mathcal{N}$, we get the same: $\mathcal{N}\subseteq\mathcal{G}$.
[…]
(vi) We have to rework the following expression
${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$ $={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}$ (60)
$={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}\,,$ (61)
where in the second line we used that ${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}){\bm{\mathrm{P}}}_{\mathcal{N}}$, and the final step is given by using that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\mathcal{N}}$. In particular:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ (62)
$=({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}))^{\top}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)\,,$ (63)
where in the last line we used that ${\bm{\mathrm{P}}}_{\mathcal{M}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{M}}$, being an orthogonal projector. ∎
B.3 Extended linear equivalence
We show that the relation defined in Definition 3 is an equivalence relation.
See 3
Proof.
To prove $\sim_{EL}$ is an equivalence relation we have to show its reflexivity, symmetry, and transitivity.
Reflexivity. Take:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\,.$ (64)
This means that there must exist ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}$ of rank $k:=\mathrm{dim}(\mathcal{M})$ such that,
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})&={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)&={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (65)
which are given by ${\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\mathcal{M}}$, ${\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\mathcal{N}}$.
Symmetry. We have to show that:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\iff(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\sim_{EL}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\,.$ (66)
This can be seen by showing one side of the implication ($\implies$):
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})&={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)&={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}\implies\begin{cases}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})&=\tilde{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)&=\tilde{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (67)
Take:
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})&={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)&={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (68)
and consider the pseudo-inverses ${\bm{\mathrm{M}}}^{+}$ and ${\bm{\mathrm{N}}}^{+}$ obtaining:
$\begin{cases}{\bm{\mathrm{M}}}^{+}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})&={\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)&={\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (69)
Notice that ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$, which follows by the fact that $\mathrm{ker}({\bm{\mathrm{M}}})^{\perp}=\tilde{}\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{N}}})^{\perp}=\tilde{}\mathcal{N}$ [60, Page 211]. Using this we have ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$, and similarly ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$. Therefore, on the right-hand side only the projectors ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$ remain, whereas we have to set $\tilde{{\bm{\mathrm{M}}}}={\bm{\mathrm{M}}}^{+}$ and $\tilde{{\bm{\mathrm{N}}}}={\bm{\mathrm{N}}}^{+}$. Both $\tilde{\bm{\mathrm{M}}}$ and $\tilde{\bm{\mathrm{N}}}$ have range $k$, as a consequence of being pseudo-inverses. Therefore, we get $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\sim_{EL}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$.
[…]
Transitivity. We have to show that:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\;\land\;(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})\sim_{EL}({\bm{\mathrm{f}}}^{*},{\bm{\mathrm{g}}}^{*})\implies({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}({\bm{\mathrm{f}}}^{*},{\bm{\mathrm{g}}}^{*})\,.$ (70)
This can be verified by substitution:
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (71)
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}\tilde{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\mathcal{M}^{*}}{{\bm{\mathrm{f}}}^{*}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)={\bm{\mathrm{N}}}\tilde{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\mathcal{N}^{*}}{{\bm{\mathrm{g}}}^{*}}_{0}(y)\end{cases}$ (72)
and by setting $\overline{{\bm{\mathrm{M}}}}={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{M}}}}$ and $\overline{{\bm{\mathrm{N}}}}={\bm{\mathrm{N}}}\tilde{{\bm{\mathrm{N}}}}$, it holds:
$\begin{cases}{\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})=\overline{{\bm{\mathrm{M}}}}{\bm{\mathrm{P}}}_{\mathcal{M}^{*}}{{\bm{\mathrm{f}}}^{*}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{P}}}_{\mathcal{N}}{{\bm{\mathrm{g}}}}_{0}(y)=\overline{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\mathcal{N}^{*}}{{\bm{\mathrm{g}}}^{*}}_{0}(y)\end{cases}$ (73)
Notice that, since ${\bm{\mathrm{M}}}:\tilde{\mathcal{M}}\to{\mathcal{M}}$ is a linear invertible transformation, and similarly $\tilde{\bm{\mathrm{M}}}:\mathcal{M}^{*}\to\tilde{\mathcal{M}}$ is also a linear invertible transformation, the composition $\overline{{\bm{\mathrm{M}}}}={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{M}}}}$ is a linear invertible transformation from $\mathcal{M}^{*}$ to $\mathcal{M}$, with $\mathrm{rank}(\overline{{\bm{\mathrm{M}}}})=k$. A similar observation also applies to $\overline{{\bm{\mathrm{N}}}}:\mathcal{N}^{*}\to\mathcal{N}$. Therefore, we have shown that:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}({\bm{\mathrm{f}}}^{*},{\bm{\mathrm{g}}}^{*})\,.$ (74)
This concludes the proof. ∎
Explicit form on ${\bm{\mathrm{N}}}$.
Based on the requirement that matrices ${\bm{\mathrm{M}}},{\bm{\mathrm{N}}}\in\mathbb{R}^{d\times\tilde{d}}$ obey ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$, we have that:
${\bm{\mathrm{N}}}:=({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}({\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}})$ (75)
Proof.
Notice that the following holds:
  • ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{M}}}^{\top}$, by the fact that ${\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}$ by Lemma 2 (v), and
  • ${\bm{\mathrm{P}}}_{\mathcal{G}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}$, since we have $({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}){\bm{\mathrm{P}}}_{\mathcal{G}}=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})$.
We use these identities to obtain:
${\bm{\mathrm{M}}}^{\top}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})$ $={\bm{\mathrm{M}}}^{\top}{{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})$ (76)
$={\bm{\mathrm{M}}}^{\top}{{\bm{\mathrm{P}}}_{\mathcal{M}}}({\bm{\mathrm{M}}}^{\top})^{+}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})$ (Using Lemma 2 (ii))
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{M}}}^{\top}}({\bm{\mathrm{M}}}^{\top})^{+}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})$ (Using ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{M}}}^{\top}$)
$={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})$ (${\bm{\mathrm{M}}}^{\top}({\bm{\mathrm{M}}}^{\top})^{+}={\bm{\mathrm{P}}}_{\mathrm{ker}({\bm{\mathrm{M}}})^{\perp}}={\bm{\mathrm{P}}}_{\mathcal{M}}$)
$={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\,.$ (77)
where we used that ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$ by Lemma 2 (iii). Notice that by Lemma 2 (vi) we have that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$. ∎
B.4 Proof of Proposition 4
See 4
Proof.
Starting from expression of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ given by the $\sim_{EL}$ equivalence relation (Definition 3):
${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ (78)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (79)
we use also the condition that ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$. We get
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$
where we used the idempotency of projectors, i.e., that ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$. ∎
B.5 Proof of Theorem 5
We begin with the following lemma, which will be used in the proof of Theorem 5.
Lemma 17.
Let $\mathcal{F}$ be a subspace of $\mathbb{R}^{d}$ and $\mathcal{M}\subseteq\mathcal{F}$ a subspace of $\mathcal{F}$. Consider $D$ elements ${\bm{\mathrm{v}}}_{i}\in\mathbb{R}^{d}$, such that the matrix
${\bm{\mathrm{F}}}:=\begin{pmatrix}{\bm{\mathrm{v}}}_{1},\ldots,{\bm{\mathrm{v}}}_{D}\end{pmatrix}$ (80)
has range $\mathrm{Im}({\bm{\mathrm{F}}})=\mathcal{F}$. Then, it holds $\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})=\mathrm{dim}(\mathcal{M})$.
Proof.
For two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [60, page 62], it follows that:
$\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})=\mathrm{rank}({\bm{\mathrm{B}}})-\mathrm{dim}(\mathrm{ker}({\bm{\mathrm{A}}})\cap\mathrm{range}({\bm{\mathrm{B}}}))\,.$ (81)
We use this formula to obtain:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})$ $=\mathrm{rank}({\bm{\mathrm{F}}})-\mathrm{dim}(\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{M}})\cap\mathrm{range}({\bm{\mathrm{F}}}))$ (82)
$=\mathrm{dim}(\mathcal{F})-\mathrm{dim}\big{(}\mathcal{M}^{\perp}\cap\mathcal{F}\big{)}$ (83)
$=\mathrm{dim}(\mathcal{F})-\mathrm{dim}\big{(}\mathbb{R}^{d}\setminus\mathcal{M}\cap\mathcal{F}\big{)}$ (84)
$=\mathrm{dim}(\mathcal{F})-\mathrm{dim}\big{(}\mathcal{F}\setminus\mathcal{M}\big{)}$ (85)
$=\mathrm{dim}(\mathcal{F})-\mathrm{dim}(\mathcal{F})+\mathrm{dim}(\mathcal{M})$ (86)
$=\mathrm{dim}(\mathcal{M})$ (87)
where the equality on the second-last line follows from the fact that $\mathcal{M}\subseteq\mathcal{F}$. ∎
We now proceed to prove Theorem 5.
See 5
Proof sketch. To prove the implication, we divide into five parts:
  1. 1.
    Starting from log-equality of probabilities, we adopt a pivoting strategy to get rid of normalizing constants [^5] A similar “pivoting” strategy is the starting point of several identifiability proofs in nonlinear ICA with auxiliary variables, e.g., [43, 24, 20, 3].;
  2. 2.
    We derive an explicit expression of ${\bm{\mathrm{M}}}$ and $\tilde{{\bm{\mathrm{M}}}}$ such that
    ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$
    ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ $=\tilde{{\bm{\mathrm{M}}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$
    hold for every ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$. We achieve that by considering $q$ points, or tokens, of $\mathcal{A}$, such that the matrices ${\bm{\mathrm{G}}}:=({\bm{\mathrm{g}}}_{0}(y_{1}),\ldots,{\bm{\mathrm{g}}}_{0}(y_{q}))$ and $\tilde{\bm{\mathrm{G}}}:=(\tilde{{\bm{\mathrm{g}}}}_{0}(y_{1}),\ldots,\tilde{{\bm{\mathrm{g}}}}(y_{q}))$ have images corresponding to $\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ and $\mathrm{SIm}(\tilde{{\bm{\mathrm{g}}}}_{0})$, respectively;
  3. 3.
    From the linear relation obtained between ${\bm{\mathrm{f}}}$ and $\tilde{\bm{\mathrm{f}}}$, we show that, having $\mathrm{dim}(\mathcal{M})=k$, also $\mathrm{dim}(\tilde{}\mathcal{M})=k$ and $\mathrm{rank}({\bm{\mathrm{M}}})=k$. By considering $\ell$ points, or sequences, $x_{i}\in\mathsf{Seq}(\mathcal{A})$, such that the matrices ${\bm{\mathrm{F}}}:=({\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{q}),\ldots,{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell}))$ and $\tilde{\bm{\mathrm{F}}}:=(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{q}),\ldots,\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell}))$ have images corresponding to $\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}})$, respectively, we will obtain that
    $k\leq m\leq k,\quad k\leq k^{\prime}\leq k\,;$
  4. 4.
    We derive an explicit expression of ${\bm{\mathrm{N}}}$ such that
    ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)$
    hold for every $y\in\mathcal{A}$. We achieve that by considering again $\ell$ points such that ${\bm{\mathrm{F}}}$ and $\tilde{{\bm{\mathrm{F}}}}$ have images corresponding to $\mathrm{SIm}({\bm{\mathrm{f}}})$ and $\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}})$, and using the linear relation obtained between ${\bm{\mathrm{f}}}$ and $\tilde{{\bm{\mathrm{f}}}}$. This allows us to derive an expression of ${\bm{\mathrm{N}}}$ that depends on ${\bm{\mathrm{M}}}$;
  5. 5.
    Finally, we show that it follows that $\mathrm{rank}({\bm{\mathrm{N}}})$ is equal to $k:=\mathrm{dim}(\mathcal{M})$. To achieve that we use the relation between ${\bm{\mathrm{g}}}_{0}$ and $\tilde{{\bm{\mathrm{g}}}}_{0}$ to show that:
    $k\leq\mathrm{rank}({\bm{\mathrm{N}}})\leq k\,.$
Recall that we are indicating with:
$\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}}),\quad\tilde{\mathcal{F}}:=\mathrm{SIm}(\tilde{{\bm{\mathrm{f}}}}),\quad\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0}),\quad\tilde{\mathcal{G}}:=\mathrm{SIm}(\tilde{{\bm{\mathrm{g}}}}_{0})\,,$
and with:
$\mathcal{M}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}),\quad\tilde{}\mathcal{M}:=\mathrm{Im}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}),\quad\mathcal{N}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{\perp},\quad\tilde{}\mathcal{N}:=\mathrm{ker}({\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}})^{\perp}\ .$ (88)
Proof.
We start by proving the implication ($\implies$).
Step 1: Using pivoting to get rid of normalizing constants. We start from the equality between log probabilities:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y)-\log Z({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}\tilde{{\bm{\mathrm{g}}}}(y)-\log\tilde{Z}({\bm{\mathrm{x}}})$ (89)
Subtracting the pivot $y_{0}\in\mathcal{A}$ to all remaining elements $y\in\mathcal{A}$, and defining ${\bm{\mathrm{g}}}_{0}:={\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0})$ and $\tilde{\bm{\mathrm{g}}}_{0}(y):=\tilde{\bm{\mathrm{g}}}(y)-\tilde{\bm{\mathrm{g}}}(y_{0})$, we get rid of the terms containing the log of the normalizing constant, i.e.,
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}({\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0}))$ $=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}(\tilde{{\bm{\mathrm{g}}}}(y)-\tilde{{\bm{\mathrm{g}}}}(y_{0}))$ (Subtract by ${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})$ and by $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}\tilde{{\bm{\mathrm{g}}}}(y_{0})$)
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\,.$ (90)
Step 2: Obtaining the relation between embeddings ${\bm{\mathrm{f}}}$ and $\tilde{\bm{\mathrm{f}}}$. We will now show that ${\bm{\mathrm{M}}}:=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}({\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}})$ and $\tilde{{\bm{\mathrm{M}}}}:=({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})^{+}(\tilde{\bm{\mathrm{G}}}^{\top})^{+}{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{{\mathcal{G}}}{\bm{\mathrm{P}}}_{{\mathcal{F}}}$ satisfies 9, showing that ${\bm{\mathrm{f}}}$ and $\tilde{{\bm{\mathrm{f}}}}$ are linearly related. For this, consider $q$ elements, or tokens, $y\in\mathcal{A}$ such that the matrix
${\bm{\mathrm{G}}}=\begin{pmatrix}{\bm{\mathrm{g}}}_{0}(y_{1}),&\cdots,&{\bm{\mathrm{g}}}_{0}(y_{q})\end{pmatrix},\quad\tilde{\bm{\mathrm{G}}}=\begin{pmatrix}\tilde{\bm{\mathrm{g}}}_{0}(y_{1}),&\cdots,&\tilde{\bm{\mathrm{g}}}_{0}(y_{q})\end{pmatrix}$ (91)
span $\mathcal{G}$ and $\tilde{\mathcal{G}}$, respectively, i.e., $\mathrm{Im}({\bm{\mathrm{G}}})=\mathcal{G}$ and $\mathrm{Im}(\tilde{\bm{\mathrm{G}}})=\tilde{}\mathcal{G}$. Taking the transpose of Equation (90) and considering these $q$ elements, we can write
${\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{G}}}}^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,.$ (92)
Next, we consider the projectors on the subspace where all functions’ images are contained. Indicate with ${\bm{\mathrm{P}}}_{\mathcal{F}}$ and ${\bm{\mathrm{P}}}_{\mathcal{G}}$ the orthogonal projectors on $\mathcal{F}$ and $\mathcal{G}$, respectively, and with ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$ the orthogonal projectors on $\tilde{\mathcal{F}}$ and $\tilde{\mathcal{G}}$, respectively. It holds
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}),\quad{\bm{\mathrm{G}}}={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{G}}},\quad\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}),\quad\tilde{\bm{\mathrm{G}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{G}}}\,,$ (93)
which can be inserted in Equation (92), leading to
${\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,.$ (94)
We make use of the pseudo-inverse of ${\bm{\mathrm{G}}}^{\top}$ to obtain:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{G}}}^{\top})^{+}}{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{G}}}^{\top})^{+}}\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (Multiply on the left by $({\bm{\mathrm{G}}}^{\top})^{+}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (From 30, $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{G}}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{A}}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,,$ (95)
where in the last line we used ${\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{G}}$ since $\mathrm{Im}({\bm{\mathrm{G}}})=\mathcal{G}$, and we denoted with ${\bm{\mathrm{A}}}:=({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}$.
Next, we insert the projectors ${\bm{\mathrm{P}}}_{\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ in Equation (95), using Lemma 2 (iii) we get:
${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{M}}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\bm{\mathrm{A}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,.$ (96)
We now consider the left pseudo-inverse of ${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}$ to obtain
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}}{\bm{\mathrm{A}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (Multiply on the left by $({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{M}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}{\bm{\mathrm{A}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (By Lemma 2 (ii) ${\bm{\mathrm{P}}}_{\mathcal{M}}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{M}}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{M}}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,,$ (97)
where we used the idempotency of the ${\bm{\mathrm{P}}}_{\mathcal{M}}$ projector, and we defined
${{\bm{\mathrm{M}}}}:=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\,.$ (98)
Following similar steps but starting from $\tilde{{\bm{\mathrm{G}}}}^{\top}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$, we arrive at a similar expression for ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}$:
${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{M}}}}{\bm{\mathrm{P}}}_{{\mathcal{M}}}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\,,$ (99)
where $\tilde{{\bm{\mathrm{M}}}}:=({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})^{+}(\tilde{\bm{\mathrm{G}}}^{\top})^{+}{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{{\mathcal{G}}}{\bm{\mathrm{P}}}_{{\mathcal{F}}}$.
Step 3: Showing that $\mathrm{dim}(\tilde{}\mathcal{M})=\mathrm{rank}({\bm{\mathrm{M}}})=\mathrm{dim}(\mathcal{M})$. Let $k:=\mathrm{dim}(\mathcal{M})$ and $k^{\prime}:=\mathrm{dim}(\tilde{\mathcal{M}})$. Also, let $m:=\mathrm{rank}({\bm{\mathrm{M}}})$. We want to show that $k^{\prime}=m=k$ and to this end we will obtain that $k\leq m\leq k$ and that $k^{\prime}\leq k\leq k$. This is done in three points:
  • (I)
    We show that $m\leq k$ from the definition of ${\bm{\mathrm{M}}}$ in Equation 98;
  • (II)
    We show that necessarily $m\geq k$ and $k^{\prime}\geq k$ from Equation 97;
  • (III)
    We show that $k^{\prime}\leq k$ from Equation 99.
(I) By equation (98), we have that ${\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{M}}}$, since by 98, in ${\bm{\mathrm{M}}}$ we have on the left ${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}$ by Lemma 2 (ii). Taking the rank of ${\bm{\mathrm{M}}}$ and using the fact that $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$ [60], we obtain:
$\mathrm{rank}({\bm{\mathrm{M}}})$ $=\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{M}}})$ (Take the rank of ${\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{M}}}$)
$m$ $\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}),\mathrm{rank}({\bm{\mathrm{M}}})\big{)}$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$m$ $\leq\min\big{(}m,k\big{)}$ (100)
$\implies m$ $\leq k$ (101)
Next, we consider $\ell$ sequences ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that the matrices:
${\bm{\mathrm{F}}}=\begin{pmatrix}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1}),&\cdots,&{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell})\end{pmatrix},\quad\tilde{\bm{\mathrm{F}}}=\begin{pmatrix}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1}),&\cdots,&\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell})\end{pmatrix}$ (102)
span the whole $\mathcal{F}$ and $\tilde{\mathcal{F}}$, respectively, i.e., $\mathrm{Im}({\bm{\mathrm{F}}})=\mathcal{F}$ and $\mathrm{Im}(\tilde{\bm{\mathrm{F}}})=\tilde{\mathcal{F}}$. Moreover, since $\mathcal{M}\subseteq\mathcal{F}$ we have that by Lemma 17 that $\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})=k$. Similarly, we have $\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{F}}})=k^{\prime}$.
(II) We consider from the Equation (97) the condition for $\ell$ points:
${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}}$ $={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{F}}}}\,.$ (103)
We evaluate the rank from Equation (103) to obtain:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})$ $=\mathrm{rank}({\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{F}}}})$ (104)
$k$ $\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{M}}}),\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{F}}}})\big{)}$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$k$ $\leq\min\big{(}m,k^{\prime}\big{)}$ (105)
$\implies m$ $\geq k,\quad k^{\prime}\geq k$ (106)
Together with 101, it shows that $k\leq m\leq k$, so it must be that $m=k$ and so $\mathrm{rank}({\bm{\mathrm{M}}})=\mathrm{dim}(\mathcal{M})$.
(III) From the Equation (99), we get similarly:
${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{F}}}}=\tilde{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}}\,.$ (107)
Following a similar proof to (II), we evaluate the rank from Equation (107) to obtain:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{F}}})$ $\leq\min\big{(}\mathrm{rank}(\tilde{\bm{\mathrm{M}}}),\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{M}}{{\bm{\mathrm{F}}}})\big{)}$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$\implies k^{\prime}$ $\leq k$ (108)
which, together with 106, it holds only as long as $k^{\prime}\leq k$. This shows that $k\leq k^{\prime}\leq k$, so it must be that $k^{\prime}=k$.
Hence, we have shown that $\mathrm{dim}(\tilde{}\mathcal{M})=\mathrm{dim}(\mathcal{M})=\mathrm{rank}({\bm{\mathrm{M}}})$. In particular, it holds that ${\bm{\mathrm{M}}}$ is an invertible map from $\tilde{}\mathcal{M}$ to $\mathcal{M}$, with pseudo-inverse ${\bm{\mathrm{M}}}^{+}$, such that:
${\mathrm{Im}({\bm{\mathrm{M}}})}=\mathcal{M},\quad{\mathrm{ker}({\bm{\mathrm{M}}})^{\perp}}={\tilde{}\mathcal{M}}\,.$ (109)
Step 4. Obtaining the relation between unembeddings ${\bm{\mathrm{g}}}$ and $\tilde{\bm{\mathrm{g}}}$. We will now show that ${\bm{\mathrm{N}}}:=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$, using the matrix ${\bm{\mathrm{M}}}$ in 98 from Step 2 satisfies Equation 10.
Similar to Step 2, we take $\ell$ points, or sequences, ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that ${\bm{\mathrm{F}}}$ and $\tilde{{\bm{\mathrm{F}}}}$ in Equation 102 span $\mathcal{F}$ and $\tilde{\mathcal{F}}$, respectively. We then have:
${\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Considering $\ell$ points for ${\bm{\mathrm{F}}}$ and $\tilde{{\bm{\mathrm{F}}}}$)
${\bm{\mathrm{F}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Using orthogonal projectors 93)
${\bm{\mathrm{F}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{M}}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Using Lemma 2 (iii))
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})^{\top}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Taking the transpose ${\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}=({\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}})^{\top}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{F}}})^{\top}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (110)
where in the last line we substituted the expression for ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}}$ given by 98. Thus, restarting from (110), and reworking the expression we get:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tilde{\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Expanding the transpose on the left)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\tilde{\bm{\mathrm{F}}}^{\top})^{+}}{\tilde{\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\tilde{\bm{\mathrm{F}}}^{\top})^{+}}\tilde{{\bm{\mathrm{F}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Multiply by pseudo-inverse $(\tilde{{\bm{\mathrm{F}}}}^{\top})^{+}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (From 30 we get $(\tilde{{\bm{\mathrm{F}}}}^{\top})^{+}(\tilde{{\bm{\mathrm{F}}}}^{\top})={\bm{\mathrm{P}}}_{\mathrm{Im}(\tilde{{\bm{\mathrm{F}}}})}={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{M}}}^{\top}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$ (111)
where we used in the last line that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$, and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$ follows by the definition of ${\bm{\mathrm{M}}}$ 98, containing on the right $({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})$. Recall that by Lemma 2 (iii), it holds that
${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}},\quad{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\,.$ (112)
Using this in 111, we obtain
${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{N}}}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (113)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{M}}}^{\top})^{+}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{M}}}^{\top})^{+}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Multiply on the left for the pseudo-inverse $({\bm{\mathrm{M}}}^{\top})^{+}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{M}}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $=({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$ (114)
where in the last line we use $({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{M}}})}={\bm{\mathrm{P}}}_{\mathcal{M}}$, where the first equality follows by 30 and the second one is given by $\mathrm{Im}({\bm{\mathrm{M}}})=\mathcal{M}$ from 109. We now use that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$ to obtain:
${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $=({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (115)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Multiply by pseudo-inverse $({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}$)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{N}}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ ($({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{N}}$, by Lemma 2 (ii))
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{N}}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{N}}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$ (116)
where we set ${\bm{\mathrm{N}}}:=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$. This expression is in line with that of Definition 3, showing the equivalence relation.
Step 5. Showing that $\mathrm{rank}({\bm{\mathrm{N}}})=\mathrm{dim}(\mathcal{N})$. It remains to show that ${\bm{\mathrm{N}}}$ has rank equal to $k:=\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathcal{M})$. Similarly to Step 3, we will show that $k\leq\mathrm{rank}({\bm{\mathrm{N}}})\leq k$ and we proceed with two points:
  • (I)
    We show that by the form of ${\bm{\mathrm{N}}}$, we obtain $\mathrm{rank}({\bm{\mathrm{N}}})\leq k$;
  • (II)
    We use 116 to obtain that $\mathrm{rank}({\bm{\mathrm{N}}})\geq k$.
(I) Notice that ${\bm{\mathrm{N}}}$ is left-invariant by multiplication to ${\bm{\mathrm{P}}}_{\mathcal{N}}$, because of the term $({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}={\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}$, by Lemma 2 (iii). Therefore:
${\bm{\mathrm{N}}}$ $={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{N}}}$ (117)
$\mathrm{rank}({\bm{\mathrm{N}}})$ $=\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{N}}})$ (118)
$\mathrm{rank}({\bm{\mathrm{N}}})$ $\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}),\mathrm{rank}({\bm{\mathrm{N}}})\big{)}$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$\mathrm{rank}({\bm{\mathrm{N}}})$ $\leq\min\big{(}k,\mathrm{rank}({\bm{\mathrm{N}}})\big{)}$ (119)
$\implies\mathrm{rank}({\bm{\mathrm{N}}})$ $\leq k$ (120)
(II) Next, consider $q$ elements of $\mathcal{A}$ such that the matrices ${\bm{\mathrm{G}}}$ and $\tilde{{\bm{\mathrm{G}}}}$ in Equation 91 have rank equal to $\mathrm{dim}(G)$ and $\mathrm{dim}(\tilde{\mathcal{G}})$, respectively. From Equation 116, it holds:
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{G}}}={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{\bm{\mathrm{G}}}$ (121)
Notice that $\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{G}}})=\mathrm{dim}(\mathcal{N})$ and $\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{{\bm{\mathrm{G}}}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 17, and it also holds $\mathrm{dim}(\mathcal{N})=\mathrm{dim}(\mathcal{M})=\mathrm{dim}(\tilde{\mathcal{M}})=\mathrm{dim}(\tilde{\mathcal{N}})$ by Lemma 2 (i) and Step 3. Let $k:=\mathrm{dim}(\mathcal{N})$. Using this in 121, we obtain:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{G}}})$ $=\mathrm{rank}({\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{\bm{\mathrm{G}}})$ (122)
$k$ $\leq\mathrm{min}\big{(}\mathrm{rank}({\bm{\mathrm{N}}}),\mathrm{rank}({\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}\tilde{\bm{\mathrm{G}}})\big{)}$ (123)
$k$ $\leq\mathrm{min}\big{(}\mathrm{rank}({\bm{\mathrm{N}}}),k)$ (124)
$\implies\mathrm{rank}({\bm{\mathrm{N}}})$ $\geq k$ (125)
This shows that, combined with 120 we have $k\leq\mathrm{rank}({\bm{\mathrm{N}}})\leq k$, which means that $\mathrm{rank}({\bm{\mathrm{N}}})=k$. Taking Steps 2, 3, 4, 5 together, we have that:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\,.$ (126)
This shows the implication.
[…]
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{EL}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\implies{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (127)
for all ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ and all $y\in\mathcal{A}$. We continue from the right-hand side to obtain:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y)-{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})$ $=\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y)-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})$ (Use explicit expression for ${\bm{\mathrm{g}}}_{0}$ and $\tilde{{\bm{\mathrm{g}}}}_{0}$)
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y)$ $=\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y)+{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})$ (Reordering all $y_{0}$ terms on the right)
$\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))$ $=\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))\cdot\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})\big{)}$ (Taking the exponential on both sides)
$\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))}{Z({\bm{\mathrm{x}}})}$ $={\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}\cdot\frac{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})\big{)}}{Z({\bm{\mathrm{x}}})}$ (Dividing by the normalizing constant $Z({\bm{\mathrm{x}}})$)
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ $={\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}\cdot\frac{1}{\tilde{Z}({\bm{\mathrm{x}}})}\,,$ (128)
where in the last line we included the expression for the conditional probability $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})=\exp({\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y))/Z({\bm{\mathrm{x}}})$ from 1, and we denoted $\tilde{Z}({\bm{\mathrm{x}}}):={Z}({\bm{\mathrm{x}}})\big{/}{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}(y_{0})-\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y_{0})\big{)}}$. To obtain the value of $\tilde{Z}$ we consider the sum over all $y\in\mathcal{A}$ for 128, giving:
$\sum_{y\in\mathcal{A}}p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ $=\sum_{y\in\mathcal{A}}{\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}\cdot\frac{1}{\tilde{Z}({\bm{\mathrm{x}}})}$ (129)
11 11 $=\sum_{y\in\mathcal{A}}{\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}\cdot\frac{1}{\tilde{Z}({\bm{\mathrm{x}}})}$ (130)
$\tilde{Z}({\bm{\mathrm{x}}})$ $=\sum_{y\in\mathcal{A}}{\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}$ (131)
which means that, from 128 we have:
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ $=\frac{\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}{\sum_{y\in\mathcal{A}}\exp(\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}(y))}$ (132)
$=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}(y\mid{\bm{\mathrm{x}}})$ (133)
showing the claim. This concludes the proof. ∎
B.6 Proof of Corollary 6
The following corollary constitutes a special case of Theorem 5, which can be easily proven by setting $d=\tilde{d}$ and requiring that $\mathcal{M}=\mathcal{N}=\mathbb{R}^{d}$.Here, we provide an alternative proof expanding previous results by Roeder et al. [3], relaxing two assumptions that were used in that context. For comparison, we report the statement by Roeder et al. [3]. To this end, fix a pivot ${\bm{\mathrm{x}}}_{0}\in\mathsf{Seq}(\mathcal{A})$ and indicate with
${\bm{\mathrm{f}}}_{0}({\bm{\mathrm{x}}}):={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})-{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{0})$ (134)
the difference between embeddings and the pivot.
Theorem (Roeder et al. [3]).
Given two models $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}}),({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$, under the assumption that:
  1. 1.
    $\mathrm{SIm}({\bm{\mathrm{f}}}_{0})=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$;
  2. 2.
    $\mathrm{SIm}(\tilde{\bm{\mathrm{f}}}_{0})=\mathrm{SIm}(\tilde{\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$;
it holds:
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}=p_{\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}}}\implies({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{L}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ (135)
where the linear equivalence relation is given by:
$({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\sim_{L}(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\iff\begin{cases}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})&={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})\\ {\bm{\mathrm{g}}}_{0}(y)&={\bm{\mathrm{N}}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\end{cases}$ (136)
$\forall{\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$ and $\forall y\in\mathcal{A}$, where ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{I}}}$ and in particular ${\bm{\mathrm{N}}}={\bm{\mathrm{M}}}^{-\top}$.
To highlight deviations, we present a proof that follows a somewhat analogous argument to the proof of Roeder et al. [3]; a direct proof may show ${\bm{\mathrm{N}}}={\bm{\mathrm{M}}}^{-1}$ relying on Theorem 5. We relax condition 2 and use the fact the assumption that $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{d}$, which is a milder condition to requiring that $\mathrm{SIm}({\bm{\mathrm{f}}}_{0})=\mathbb{R}^{d}$. We prove the following:
See 6
Proof.
Our proof follows a similar technique to Lachapelle et al. [25, Theorem B.4].
Identifiability of ${\bm{\mathrm{f}}}$. First notice that from the equivalnce of log-likelihood we can write:
$\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ $=\log p_{\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{x}}})$ (137)
${\bm{\mathrm{g}}}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})-\log Z({\bm{\mathrm{x}}})$ $=\tilde{{\bm{\mathrm{g}}}}(y)^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})-\log\tilde{Z}({\bm{\mathrm{x}}})$ (138)
$({\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=(\tilde{{\bm{\mathrm{g}}}}(y)-\tilde{{\bm{\mathrm{g}}}}(y_{0}))^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (139)
${\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=\tilde{{\bm{\mathrm{g}}}}_{0}(y)^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (140)
where in the last line we subtracted the pivot for different log-probabilities on the points $y$ and $y_{0}$. We now consider the matrix ${\bm{\mathrm{G}}}$ constructed to contain $d$ differences:
${\bm{\mathrm{G}}}=\begin{pmatrix}{\bm{\mathrm{g}}}_{0}(y_{1}),&\cdots&{\bm{\mathrm{g}}}_{0}(y_{d})\end{pmatrix}$ (141)
such that it is invertible. Since by the diversity condition $\mathrm{SIm}({\bm{\mathrm{g}}}_{0})=\mathbb{R}^{d}$, such a marix always exists. Let $\tilde{{\bm{\mathrm{G}}}}$ be the corresponding matrix of differences for $\tilde{{\bm{\mathrm{g}}}}$, we obtain:
${\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=\tilde{{\bm{\mathrm{G}}}}^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (142)
Then, we obtain:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{G}}}^{-\top}}\tilde{{\bm{\mathrm{G}}}}^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (143)
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{M}}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (144)
where we denoted as ${\bm{\mathrm{M}}}={\bm{\mathrm{G}}}^{-\top}\tilde{{\bm{\mathrm{G}}}}^{\top}$. Next, since $\mathrm{SIm}({\bm{\mathrm{f}}})=\mathbb{R}^{d}$ we can consider $d$ elements ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that
${\bm{\mathrm{F}}}=\begin{pmatrix}{\bm{\mathrm{f}}}(x_{1}),&\cdots,&{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{d})\end{pmatrix}$ (145)
is invertible. Let $\tilde{{\bm{\mathrm{F}}}}$ be the corresponding matrix for $\tilde{{\bm{\mathrm{f}}}}$. In this way we obtain the following:
${\bm{\mathrm{F}}}={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{F}}}}\,,$ (146)
and since ${\bm{\mathrm{F}}}$ is invertible, it must be that also ${\bm{\mathrm{M}}}$ and ${\bm{\mathrm{F}}}$ are invertible matrices of rank $d$. This shows that
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (147)
${\bm{\mathrm{M}}}\in\mathbb{R}^{d\times d}$ is invertible.
Identifiability of ${\bm{\mathrm{g}}}_{0}$. Next, we consider the implication for ${\bm{\mathrm{g}}}_{0}$. We start again from the pivot difference of Equation 140:
${\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=\tilde{{\bm{\mathrm{g}}}}_{0}(y)^{\top}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (148)
${\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=\tilde{{\bm{\mathrm{g}}}}_{0}(y)^{\top}{\bm{\mathrm{M}}}^{-1}{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (149)
where we substituted $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})={\bm{\mathrm{M}}}^{-1}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ from 95. Therefore, taking $d$ points ${\bm{\mathrm{x}}}\in\mathsf{Seq}(\mathcal{A})$, such that the matrix ${\bm{\mathrm{F}}}$ is invertible, restarting from the transpose of 149 we obtain:
${\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{M}}}^{-\top}\tilde{{\bm{\mathrm{g}}}}_{0}(y)$ (150)
${\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{M}}}^{-\top}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)$ (Multiplying for the inverse of ${\bm{\mathrm{F}}}^{\top}$)
${\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{M}}}^{-\top}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\,,$ (151)
This means that we have:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})_{\phantom{0}}$ $={\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})_{\phantom{0}}$ (152)
${\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{N}}}\tilde{{\bm{\mathrm{g}}}}_{0}(y)\,,$ (153)
where we have defined ${\bm{\mathrm{N}}}:={\bm{\mathrm{M}}}^{-\top}$, such that ${\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}={\bm{\mathrm{I}}}$, proving the claim. ∎
Appendix C Additional Results and Proofs of Section 4
C.1 Relational Linear Steering Property
We here want to discuss an additional linear property besides those presented in Section 4, termed linear steering property. This behavior is also referred to as the linear intervening property by Park et al. [23].
It has been empirically observed that there exist steering vectors that influence next-token predictions [23, 2, 61, 62], in the following sense: If ${\bm{\mathrm{v}}}$ encode the average difference between English to Italian embeddings, adding ${\bm{\mathrm{v}}}$ to ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ for the sentence ${\bm{\mathrm{s}}}=$The king sits on the” would change the most-likely next token prediction $y=$throne” to $\hat{y}=$trono”, and similarly this applies for other sentences, affecting the most-likely next-token prediction to move from the English token to the Italian counterpart.
We define this property as follows:
Definition 18 (Relational Linear Steering).
We say that a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ possess linear relational steering for ${\bm{\mathrm{q}}}_{0}$ and the set of $\{{\bm{\mathrm{q}}}_{1},\ldots,{\bm{\mathrm{q}}}_{m}\}$, for $m\geq 1$ queries ${\bm{\mathrm{q}}}_{j}\neq{\bm{\mathrm{q}}}_{0}$, if (1) it linearly represents ${\bm{\mathrm{q}}}_{0}$ in $\Gamma_{0}$ and all ${\bm{\mathrm{q}}}_{j}$ on $\Gamma_{j}$, and (2) there exists a vector ${\bm{\mathrm{v}}}\in\mathbb{R}^{d}$ such that:
${\bm{\mathrm{P}}}_{\Gamma_{0}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{0}}{\bm{\mathrm{v}}}\neq\mathbf{0},\quad{\bm{\mathrm{P}}}_{\Gamma_{j}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{j}}{\bm{\mathrm{v}}}=\mathbf{0},\quad\forall j\in[m]$ (154)
We prove that relational linearity allows for this property:
Proposition 19.
If (1) $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents ${\bm{\mathrm{q}}}_{0}$ on $\Gamma_{0}$ (Definition 7) and (2) $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents $m\geq 1$ queries ${\bm{\mathrm{q}}}_{j}\neq{\bm{\mathrm{q}}}_{0}$ on $\Gamma_{j}$, such that $\big{(}\bigcup_{j}\Gamma_{{\bm{\mathrm{q}}}_{j}}\big{)}\cap\Gamma_{{\bm{\mathrm{q}}}_{0}}\subsetneq\Gamma_{{\bm{\mathrm{q}}}_{0}}$, then the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ satisfies linear relational steering for ${\bm{\mathrm{q}}}_{0}$ and the set of queries $\{{\bm{\mathrm{q}}}_{1},\ldots,{\bm{\mathrm{q}}}_{m}\}$.
Proof.
From the assumptions (1) and (2) we have that relational linearity as per Definition 7 implies that:
$\begin{cases}{\bm{\mathrm{P}}}_{\Gamma_{0}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{0})&={\bm{\mathrm{P}}}_{\Gamma_{0}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{0}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma_{0}}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}_{0}}\\ {\bm{\mathrm{P}}}_{\Gamma_{j}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{j})&={\bm{\mathrm{P}}}_{\Gamma_{j}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{j}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma_{j}}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}_{j}},\quad\forall j\in[\ell]\end{cases}$ (155)
Let $\Gamma_{0,int}=\bigcup_{j=1}^{\ell}\Gamma_{{\bm{\mathrm{q}}}_{j}}\cap\Gamma_{{\bm{\mathrm{q}}}_{0}}$. By assumption (2), $\Gamma_{0,int}\subsetneq\Gamma_{0}$, meaning that it exists a non-empty $\Gamma_{0,\perp}=\Gamma_{{\bm{\mathrm{q}}}_{0}}\setminus\Gamma_{0,int}$. Let ${\bm{\mathrm{v}}}\in\Gamma_{0,\perp}$. It holds:
${\bm{\mathrm{P}}}_{\Gamma_{0}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{0}}{\bm{\mathrm{v}}}={\bm{\mathrm{v}}}$ (156)
because $\Gamma_{0,\perp}\subseteq\Gamma_{{\bm{\mathrm{q}}}_{0}}$. Proceeding similarly, it holds $\Gamma_{0,\perp}\cap\Gamma_{{\bm{\mathrm{q}}}_{j}}=\varnothing$ $\forall j\in[\ell]$, by assumption (2). Therefore, we have:
${\bm{\mathrm{P}}}_{\Gamma_{j}}{\bm{\mathrm{A}}}_{{\bm{\mathrm{q}}}_{j}}{\bm{\mathrm{v}}}=\mathbf{0}$ (157)
showing the claim. ∎
[…]
C.2 Proof of Lemma 9
See 9
Proof.
[…]
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}})}$ $=\beta\cdot\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{2}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{3}\mid{\bm{\mathrm{s}}})}$ (158)
$\log\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}(y_{0}))}{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}(y_{1}))}$ $=\beta\cdot\log\frac{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}(y_{2}))}{\exp({\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}(y_{3}))}$ (159)
$\log{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}({\bm{\mathrm{g}}}(y_{0})-{\bm{\mathrm{g}}}(y_{1}))\big{)}}$ $=\beta\cdot\log{\exp\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}({\bm{\mathrm{g}}}(y_{2})-{\bm{\mathrm{g}}}(y_{3}))\big{)}}$ (160)
${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}({\bm{\mathrm{g}}}(y_{0})-{\bm{\mathrm{g}}}(y_{1}))$ $=\beta\cdot{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}({\bm{\mathrm{g}}}(y_{2})-{\bm{\mathrm{g}}}(y_{3}))$ (161)
And substituting ${\bm{\mathrm{g}}}_{1}(y_{0}):={\bm{\mathrm{g}}}(y_{0})-{\bm{\mathrm{g}}}(y_{1})$ and ${\bm{\mathrm{g}}}_{3}(y_{2}):={\bm{\mathrm{g}}}(y_{2})-{\bm{\mathrm{g}}}(y_{3})$ we obtain:
${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}_{1}(y_{0})=\beta\cdot{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})^{\top}{\bm{\mathrm{g}}}_{3}(y_{2})$ (162)
Consider $\ell$ elements ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$, such that:
${\bm{\mathrm{F}}}=({\bm{\mathrm{f}}}({\bm{\mathrm{s}}}_{1}),\ldots,{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}_{\ell}))$ (163)
spans $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$. We then obtain:
${\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{g}}}_{1}(y_{0})=\beta{\bm{\mathrm{F}}}^{\top}{\bm{\mathrm{g}}}_{3}(y_{2})$ (164)
and multiplying both sides of (164) from the left with the pseudo-inverse of ${\bm{\mathrm{F}}}^{\top}$ we get:
${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{g}}}_{1}(y_{0})=\beta{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{g}}}_{3}(y_{2}).$ (165)
Notice that, both ${\bm{\mathrm{g}}}_{1}(y_{0}),{\bm{\mathrm{g}}}_{3}(y_{2})\in\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$, then it holds ${\bm{\mathrm{g}}}_{1}(y_{0})={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{1}(y_{0})$ ${\bm{\mathrm{g}}}_{3}(y_{2})={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{3}(y_{2})$. Using this we obtain:
${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{1}(y_{0})$ $=\beta{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{3}(y_{2})$ (166)
$({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{1}(y_{0})$ $=\beta({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{3}(y_{2})$ (167)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{1}(y_{0})$ $=\beta{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{3}(y_{2})\,.$ (168)
Regrouping the two terms on one side we can see that to have parallelism in $\mathcal{N}$ (Definition 8), we must have the following:
${\bm{\mathrm{P}}}_{\mathcal{N}}({\bm{\mathrm{g}}}_{1}(y_{0})-\beta\cdot{\bm{\mathrm{g}}}_{3}(y_{2}))=\mathbf{0}$ (169)
i.e., ${\bm{\mathrm{g}}}_{1}(y_{0})-\beta\cdot{\bm{\mathrm{g}}}_{3}(y_{2})\in\mathcal{N}^{\perp}$. Therefore, ${\bm{\mathrm{g}}}_{1}(y_{0})$ and ${\bm{\mathrm{g}}}_{3}(y_{2})$ are parallel in $\mathcal{N}$.
The implication ($\implies$) is given by a similar proof by starting from Definition 8, i.e., that ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{1}(y_{0})=\beta{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{3}(y_{2})$. Therefore by multiplying the two for any embedding ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ we get
${\bm{\mathrm{g}}}_{1}(y_{0})^{\top}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ $=\beta{\bm{\mathrm{g}}}_{3}(y_{2})^{\top}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ (170)
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}})}$ $=\beta\cdot\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{2}\mid{\bm{\mathrm{s}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{3}\mid{\bm{\mathrm{s}}})}\,.$ (171)
This shows the claim.
Remark 20.
In this context, we exclude the case $\beta=0$ because, for the zero vector ${\bm{\mathrm{a}}}=0$, a contradiction arises. Specifically, any vector ${\bm{\mathrm{b}}}\in\mathbb{R}^{d}$ would be trivially parallel to ${\bm{\mathrm{a}}}$ (i.e., ${\bm{\mathrm{a}}}=\beta{\bm{\mathrm{b}}}$ with $\beta=0$). However, conversely, we would obtain that no scalar $\beta\in\mathbb{R}$ exists such that ${\bm{\mathrm{b}}}=\beta{\bm{\mathrm{a}}}$.
C.3 Proof of Proposition 11
See 11
Proof.
We start from the relational linearity as per Definition 7 for ${\bm{\mathrm{q}}}$ on $\Gamma\subset\mathbb{R}^{d}$:
${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,,$ (172)
and recall that $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})=\{{\bm{\mathrm{A}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{v}}}\mid{\bm{\mathrm{v}}}\in\Gamma\}$. By assumption (2) $\Gamma_{\bm{\mathrm{q}}}\subseteq\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. Then, for any pair $y_{i},y_{j}\in\mathcal{A}$ such that ${\bm{\mathrm{g}}}_{i}(y):={\bm{\mathrm{g}}}(y_{j})-{\bm{\mathrm{g}}}(y_{i})\in\Gamma_{\bm{\mathrm{q}}}$, we can find a vector ${\bm{\mathrm{\gamma}}}\in\Gamma$ such that
${\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{\gamma}}}={\bm{\mathrm{g}}}_{i}(y_{j})\,.$ (173)
Notice that from this expression we can also write:
${\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{\gamma}}}={\bm{\mathrm{g}}}_{i}(y_{j})\,,$ (174)
since ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{\gamma}}}={\bm{\mathrm{\gamma}}}$. Hence, the vector ${\bm{\mathrm{\gamma}}}$ can be obtained taking the pseudo-inverse of ${\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma}$:
$({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{\gamma}}}$ $=({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}{\bm{\mathrm{g}}}_{i}(y_{j})$ (175)
${\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}})}{\bm{\mathrm{\gamma}}}$ $=({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}{\bm{\mathrm{g}}}_{i}(y_{j})$ (176)
Notice that, since $({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}=({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}$ and ${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}{\bm{\mathrm{g}}}_{i}(y_{j})={\bm{\mathrm{g}}}_{i}(y_{j})$, we have that $({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})^{+}{\bm{\mathrm{g}}}_{i}(y_{j})=\mathbf{0}$ only when ${\bm{\mathrm{g}}}_{i}(y_{j})=\mathbf{0}$. As a consequence, when ${\bm{\mathrm{g}}}_{i}(y_{j})\neq\mathbf{0}$, we have that also ${\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}})}{\bm{\mathrm{\gamma}}}\neq\mathbf{0}$. Fix this ${\bm{\mathrm{\gamma}}}$ and consider:
${\bm{\mathrm{\gamma}}}^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ $={\bm{\mathrm{\gamma}}}^{\top}\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})-{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\big{)}$ (177)
$({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{\gamma}}})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ $={\bm{\mathrm{\gamma}}}^{\top}\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})-{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\big{)}$ (178)
${\bm{\mathrm{g}}}_{i}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ $={\bm{\mathrm{\gamma}}}^{\top}\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})-{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\big{)}\,,$ (179)
which shows that $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents $\Gamma_{\bm{\mathrm{q}}}$ related to ${\bm{\mathrm{q}}}$, showing the claim. ∎
C.4 Proof of Proposition 13
See 13
Proof.
Under the assumption (2), take a pivot $y_{j}\in\mathcal{Y}_{P}$, then for all remaining $y_{i}\in\mathcal{Y}_{P}$ denote with ${{\bm{\mathrm{g}}}_{j}(y_{i}):={\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j})\in\Gamma}$. Taking the log-ratios between the conditional probabilities
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P}),\;\text{ and }p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{j}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})\,,$ (180)
for conditional probabilities restricted to $\mathcal{Y}_{P}$, as in Definition 12, we obtain:
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{j}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ $=\log\frac{\exp\big{(}{\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\big{)}}{\exp\big{(}{\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\big{)}}$ (181)
$=\log\exp\big{(}({\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\big{)}$ (182)
$=\big{(}({\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\big{)}$ (183)
$={\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\,.$ (184)
Due to relational linearity of ${\bm{\mathrm{q}}}$ onto $\Gamma$ (Assumption (1)), we can write following (see Definition 7):
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,,$ (185)
we can then take any ${\bm{\mathrm{g}}}_{j}(y_{i})\in\Gamma$, and multiply their transpose times both sides of (185) from the right. We then get
${\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (186)
$({\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{g}}}_{j}(y_{i}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $=({\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{g}}}_{j}(y_{i}))^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+({\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{g}}}_{j}(y_{i}))^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (187)
${\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (188)
We can then substitute the expression on the RHS in (188) to (184) to obtain:
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{j}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}={\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (189)
Now take the conditional probability $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})$ which can be written as
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})$ $=\frac{e^{{\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (190)
$=\frac{e^{{\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}\frac{e^{{\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{e^{{\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}$ (Multiply and divide by the same term)
$=\frac{e^{({\bm{\mathrm{g}}}(y_{i})-{\bm{\mathrm{g}}}(y_{j}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}e^{{\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ (Rearrange terms in the exponential)
$=\frac{e^{{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}e^{{\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ (Substitute ${\bm{\mathrm{g}}}_{j}(y_{i})$)
$=\exp\big{(}{{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{g}}}_{j}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}}\big{)}\,\frac{e^{({\bm{\mathrm{g}}}(y_{j})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (Use 188)
$=\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})}\big{)}\,\exp\big{(}-{{\bm{\mathrm{g}}}(y_{j})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})}\big{)}\,\frac{e^{({\bm{\mathrm{g}}}(y_{j})\top{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (Separate the term depending on $y_{i}$ to those that do not)
$=\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}\,\frac{e^{{\bm{\mathrm{g}}}(y_{j})^{\top}\big{(}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})-{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\big{)}}}{Z({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})}$ (Rearrange exponential on the right)
$=\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}C\,,$ (191)
where we denoted with $C$ the scaling factor applied to the exponential, which we will treat as a constant since it does not depend on $y_{i}\in\mathcal{A}$. From this expression, take the sum on $\mathcal{Y}_{P}$ to obtain that:
$\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}\,C$ $=\sum_{y_{i}\in\mathcal{Y}_{P}}p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})$ (192)
$\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}\,C$ $=1$ (The sum on the right equals to $1$)
$C$ $=1\big{/}\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}$ (193)
Denote with ${\bm{\mathrm{w}}}{{}_{i}}:={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{g}}}(y_{i})$ and with $b_{i}:={\bm{\mathrm{g}}}(y_{i})^{\top}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$. Then using this and 193 inside 191 we get
$p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{i}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}};\mathcal{Y}_{P})$ $=\exp\big{(}{{\bm{\mathrm{g}}}(y_{i})^{\top}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}})\big{)}}C$ (194)
$=\frac{\exp\big{(}{\bm{\mathrm{w}}}_{i}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}{\sum_{y_{i}\in\mathcal{Y}_{P}}\exp\big{(}{\bm{\mathrm{w}}}_{i}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+b_{i}\big{)}}$ (Substitute for ${\bm{\mathrm{w}}}_{i}$ and $b_{i}$, and 193)
$=\mathrm{softmax}\big{(}{\bm{\mathrm{W}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{b}}}\big{)}_{i}\,,$ (195)
where we defined ${\bm{\mathrm{W}}}:=({\bm{\mathrm{w}}}_{1},\ldots,{\bm{\mathrm{w}}}_{\ell})^{\top}$ and ${\bm{\mathrm{b}}}:=(b_{1},\ldots,b_{\ell})^{\top}$. This shows that the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ can be linear probed for ${\bm{\mathrm{q}}}$ in $\mathcal{Y}_{P}$ with ${\bm{\mathrm{W}}}$ and ${\bm{\mathrm{b}}}$. ∎
Appendix D Proof of Section 5
D.1 Proof of Proposition 14
See 14 Proof Sketch. The proof is divided in two steps:
  1. 1.
    We first prove the implication that $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$ linearly represents ${\bm{\mathrm{q}}}$ on a subset $\tilde{\Gamma}\subseteq\mathbb{R}^{d}$;
  2. 2.
    Then we show that $\tilde{\Gamma}\subseteq\mathcal{N}$ and that $\tilde{\Gamma}_{\bm{\mathrm{q}}}:=\mathrm{Im}(\tilde{{\bm{\mathrm{A}}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma})\subseteq\tilde{}\mathcal{N}$.
Proof.
Step 1. We begin from the relational linearity definition for model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$. It holds
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,,$ (196)
where $\Gamma\subseteq\mathcal{N}$. This also means that ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}={\bm{\mathrm{P}}}_{\Gamma}$. Denote with $\mathcal{F}:=\mathrm{SIm}({\bm{\mathrm{f}}})$ and with $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$. By assumption, it holds that $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma})=\{{\bm{\mathrm{A}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{v}}}\mid{\bm{\mathrm{v}}}\in\Gamma\}$ is a subset of $\mathcal{M}$. This implies, in turn, that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{A}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma}={\bm{\mathrm{A}}}^{\top}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma}$. We use this to write:
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (Using ${\bm{\mathrm{P}}}_{\Gamma}={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}$)
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (Using ${\bm{\mathrm{f}}}={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}$ on the left and ${\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}={\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\mathcal{M}}$ on the right)
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,,$ (197)
where in the last line we used that ${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}={\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{P}}}_{\mathcal{M}}$ by Lemma 2 (vi). We now substitute the expression for $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$ based on the RHS of the equivalence relation(9) in 197 to get
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (198)
Now starting from ${\bm{\mathrm{N}}}^{\top}{\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$, as specified in Definition 3, we can apply the following steps:
${\bm{\mathrm{N}}}^{\top}{\bm{\mathrm{M}}}$ $={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ (199)
$({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{N}}}^{\top}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ (Multiply on the left by the pseudo-inverse $({\bm{\mathrm{N}}}^{\top})^{+}$)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ (Using $({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{N}}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{N}}$)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ (Idempotency of the orthogonal projector ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}$ (Substitute ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ from Lemma 2 (vi))
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}$ (Using ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$)
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}$ $=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}$ (Using $({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}=({\bm{\mathrm{N}}}^{\top})^{+}$)
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ $={\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}\,.$ (200)
where in the last line we multiplied on the left by the orthogonal projector ${\bm{\mathrm{P}}}_{\Gamma}$ and used that ${\bm{\mathrm{M}}}={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$. We can now show that we can substitute 200 into the left-hand side of 198:
${\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (201)
${\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (Use ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{\bm{\mathrm{f}}}=\tilde{\bm{\mathrm{f}}}$)
$({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})+({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (Multiply on the left by $({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}$)
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $=\tilde{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})+\tilde{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,,$ (202)
where we denoted with ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ the orthogonal projector on $\tilde{\Gamma}$ and we defined:
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ $:=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}$ (203)
$\tilde{{\bm{\mathrm{A}}}}_{\bm{\mathrm{q}}}$ $:=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}$ (204)
$\tilde{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ $:=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (205)
Multiplying by ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ we arrive at the same expression for linearity for the model $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})$:
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\tilde{\Gamma}}\tilde{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}\tilde{{\bm{\mathrm{f}}}}+{\bm{\mathrm{P}}}_{\tilde{\Gamma}}\tilde{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (206)
Step 2. We proceed to show that $\tilde{\Gamma}\subseteq\tilde{\mathcal{N}}$. First, we have to show that ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}={\bm{\mathrm{P}}}_{\tilde{\Gamma}}$. To this end, notice that $({\bm{\mathrm{N}}}^{\top})^{+}:\tilde{\mathcal{N}}\to\mathcal{N}$. Therefore, we have from Equation 203:
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ $=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}$ (207)
$=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$ (Using $({\bm{\mathrm{N}}}^{\top})^{+}=({\bm{\mathrm{N}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$)
$={\bm{\mathrm{P}}}_{\tilde{\Gamma}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\,.$ (208)
Taking the transpose of ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ we obtain:
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}^{\top}$ $=({\bm{\mathrm{P}}}_{\tilde{\Gamma}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}})^{\top}$ (209)
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ $={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\Gamma}}^{\top}$ (210)
${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ $={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ (211)
where we used that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}^{\top}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}^{\top}={\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ because both are symmetric matrices. This means, in turn, that ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$ and ${\bm{\mathrm{P}}}_{\tilde{\Gamma}}$ commute, and so $\tilde{\Gamma}$ must be a contained in $\tilde{\mathcal{N}}$. Similarly from the expression of $\tilde{{\bm{\mathrm{A}}}}_{\bm{\mathrm{q}}}$ we get:
$\tilde{{\bm{\mathrm{A}}}}_{\bm{\mathrm{q}}}$ $=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}$ (212)
$=({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{+}{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ (213)
$=\tilde{{\bm{\mathrm{A}}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\,.$ (214)
This means that $\tilde{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\Gamma}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\Gamma}}$, and so $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}(\tilde{{\bm{\mathrm{A}}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\Gamma}})\subseteq\tilde{\mathcal{M}}$. To find the expression for $\tilde{\Gamma}$, we consider the following:
$\tilde{\Gamma}$ $=\mathrm{ker}({\bm{\mathrm{P}}}_{\Gamma}({\bm{\mathrm{N}}}^{\top})^{+})^{\perp}$ (215)
$=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma}^{\top})$ ($\mathrm{ker}({\bm{\mathrm{A}}})^{\perp}=\mathrm{Im}({\bm{\mathrm{A}}}^{\top})$, for any matrix ${\bm{\mathrm{A}}}$ [60])
$=\mathrm{Im}({\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\Gamma})\,.$ (216)
This proves the claim. ∎
What happens if $\Gamma_{\bm{\mathrm{q}}}\not\subseteq\mathcal{M}$?. We discuss the case when the condition in Proposition 14 is not met due to $\Gamma_{\bm{\mathrm{q}}}:=\mathrm{Im}({\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}^{\top}{\bm{\mathrm{P}}}_{\Gamma})\not\subseteq\mathcal{M}$. We show that even if a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ linearly represents ${\bm{\mathrm{q}}}$ on $\Gamma$, a $\sim_{EL}$-equivalent model $(\tilde{{\bm{\mathrm{f}}}},\tilde{\bm{\mathrm{g}}})$ may not. This is due to the fact that the information contained in $\mathcal{F}\setminus\mathcal{M}$, used to relationally represent ${\bm{\mathrm{q}}}$, may not be linearly transformed on another $\sim_{EL}$-equivalent model.
In fact, when $\Gamma_{\bm{\mathrm{q}}}\not\subseteq\mathcal{M}$, consider the expression for relational linearity given by Definition 7 where
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (217)
We take the first term on the RHS of (217): by inserting the projector ${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}$, we rewrite it as follows:
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{f}}}({\bm{\mathrm{s}}})\,,$ (218)
where we used the identity ${\bm{\mathrm{I}}}={\bm{\mathrm{P}}}_{\mathcal{M}}+({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}})$ to separate the contributions, inside and outside $\mathcal{M}$. We can thus rewrite 217 as:
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}$ (Introduce ${\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}={\bm{\mathrm{f}}}$)
$\propto{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ (${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}{\bm{\mathrm{P}}}_{\mathcal{M}}=\mathbf{0}$)
$={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}){\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ (Multiplying $({\bm{\mathrm{I}}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}$)
$={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ (219)
where the in the last equation we used that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}={\bm{\mathrm{P}}}_{\mathcal{M}}$ by Lemma 2 (v). To show that this can lead to non-linearities implying deviations from relational linearity, suppose that:
${\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ $={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})+({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}}){\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$ (220)
$={{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})}+({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}})\tilde{{\bm{\mathrm{f}}}}^{2}({\bm{\mathrm{x}}})$ (221)
where the first term follows from the equivalence relation 9, i.e., ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})={\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}})$, and we used $\tilde{\bm{\mathrm{f}}}^{2}({\bm{\mathrm{x}}})=(\tilde{f}_{1}({\bm{\mathrm{x}}})^{2},\ldots,\tilde{f}_{\tilde{d}}({\bm{\mathrm{x}}})^{2})$ to denote the square of the components of $\tilde{{\bm{\mathrm{f}}}}$. Notice that, this choice is allowed since the components of ${\bm{\mathrm{f}}}$ outside $\mathcal{M}$, i.e., those in $\mathcal{F}\setminus\mathcal{M}$, can be arbitrarily chosen and do not contribute to the dot-product with ${\bm{\mathrm{g}}}_{0}$. Therefore, substituting this expression to 219 we get:
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\propto{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}})\tilde{{\bm{\mathrm{f}}}}^{2}({\bm{\mathrm{s}}})$ (222)
and substituting ${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{N}}}^{+}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$, implied by the LHS of 197 and from the equality in 200, we have that:
${\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{N}}}^{+}\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\propto{\bm{\mathrm{P}}}_{\Gamma}{\bm{\mathrm{A}}}_{\bm{\mathrm{q}}}{\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}})\tilde{{\bm{\mathrm{f}}}}^{2}({\bm{\mathrm{s}}})\,,$ (223)
which shows a non-linear dependence of $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ on $\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{s}}})$, invalidating relational linearity when ${\bm{\mathrm{P}}}_{\Gamma_{\bm{\mathrm{q}}}}({\bm{\mathrm{P}}}_{\mathcal{F}}-{\bm{\mathrm{P}}}_{\mathcal{M}})\neq\mathbf{0}$.
D.2 Proof of Proposition 15
See 15
Proof.
Given that ${\bm{\mathrm{\gamma}}}$ and ${\bm{\mathrm{\gamma}}}^{\prime}$ are parallel in $\mathcal{N}$ (Definition 8), we have that:
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}=\beta{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}^{\prime}\,,$ (224)
where $\beta\neq 0$ is given by $\beta={\lVert{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}\rVert}/{\lVert{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}^{\prime}\rVert}$ (see also Remark 20. We consider the components of the $\sim_{EL}$-equivalent model, given by
${\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}},\quad{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}^{\prime}={\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}\,.$ (225)
Using this in 224 we get:
${\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}=\beta{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ (226)
and multiplying from the left by the pseudoinverse of ${\bm{\mathrm{N}}}$ we get:
${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}$ $=\beta{\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ (227)
${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}$ $=\beta{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ (Using ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$)
${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}$ $=\beta{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ (228)
which shows that $\tilde{\bm{\mathrm{\gamma}}}$ is parallel to $\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ in ~$\tilde{}\mathcal{N}$. To prove the reverse implication, the same steps can be repeated by symmetry, taking:
${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}={\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}},\quad{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{\gamma}}}^{\prime}={\bm{\mathrm{N}}}^{+}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{\gamma}}}^{\prime}\,.$ (229)
for two vectors $\tilde{\bm{\mathrm{\gamma}}}$ and $\tilde{\bm{\mathrm{\gamma}}}^{\prime}$ parallel in ~$\tilde{}\mathcal{N}$, giving the desired result. ∎
Appendix E Context-query-reply sentences: Corner cases
E.1 Paraphrases
[…]
Here, we analyze how paraphrastic aspects of textual data can be described with relational context-query-reply (${\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}\mathbin{\smallfrown}y$) strings for a model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$. We start by providing a tentative definition of of praphrasatic sentences in terms of their entailed conditional log-probabilities for different pairs of next-tokens.
Definition 21 (Paraphrases).
We say that ${\bm{\mathrm{q}}}_{2}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{2}\subseteq\mathcal{A}$ is a paraphrase of ${\bm{\mathrm{q}}}_{1}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{1}\subseteq\mathcal{A}$ for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ if there exists $\beta\neq 0$ such that, for all $y_{0},y_{1}\in\mathcal{Y}_{1}\subseteq\mathcal{A}$, there exists $\hat{y}_{0},\hat{y}_{1}\in\mathcal{Y}_{2}\subseteq\mathcal{A}$, for which it holds:
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})}=\beta\cdot\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(\hat{y}_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(\hat{y}_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})}\,.$ (230)
[…]
We show that sentences and next-tokens as in Definition 21 induce a specific structure in the model embeddings, linearly relating the representations ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ and ${\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})$. To this end, we will define $\mathrm{SIm}({\bm{\mathrm{g}}}_{0})_{\mathcal{Y}}:=\mathrm{span}\{{\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0})\mid\;y_{0},y\in\mathcal{Y}\}$ and $\mathrm{SIm}({\bm{\mathrm{f}}})_{\bm{\mathrm{q}}}:=\mathrm{span}\{{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})\mid{\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})\}$.
Proposition 22.
If (1) ${\bm{\mathrm{q}}}_{2}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{2}\subseteq\mathcal{A}$ is a paraphrase of ${\bm{\mathrm{q}}}_{1}\in\mathsf{Seq}(\mathcal{A})$ on $\mathcal{Y}_{1}\subseteq\mathcal{A}$ for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ and (2) for the subspaces $\Gamma_{1}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})_{\mathcal{Y}_{1}}$ and $\Gamma_{2}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})_{\mathcal{Y}_{2}}$, it holds that $\Gamma_{1}\subseteq\mathrm{SIm}({\bm{\mathrm{f}}})_{{\bm{\mathrm{q}}}_{1}}=:\mathcal{F}_{1}$ and $\Gamma_{2}\subseteq\mathrm{SIm}({\bm{\mathrm{f}}})_{{\bm{\mathrm{q}}}_{2}}=:\mathcal{F}_{2}$, then $\mathrm{dim}(\Gamma_{1})=\mathrm{dim}(\Gamma_{2})$ and there exists a matrix ${\bm{\mathrm{O}}}\in\mathbb{R}^{d\times d}$ that defines a linear, invertible transformation from $\Gamma_{2}$ to $\Gamma_{1}$ such that
${\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})=\beta{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\,.$ (231)
Proof.
We start with the equality between logs of probabilities given by Definition 21:
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})}$ $=\beta\cdot\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(\hat{y}_{1}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(\hat{y}_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})}$ (232)
$\log\exp\big{(}({\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})\big{)}$ $=\beta\log\exp\big{(}({\bm{\mathrm{g}}}(\hat{y}_{1})-{\bm{\mathrm{g}}}(\hat{y}_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\big{)}$ (Substituting 1 for the conditional probabilities)
$({\bm{\mathrm{g}}}(y_{1})-{\bm{\mathrm{g}}}(y_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ $=\beta({\bm{\mathrm{g}}}(\hat{y}_{1})-{\bm{\mathrm{g}}}(\hat{y}_{0}))^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\,,$ (233)
where in the last line we removed the logarithm of the exponential on both sides. Define ${\bm{\mathrm{g}}}_{0}(y):={\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0})$ for a pivot $y_{0}\in\mathcal{Y}_{1}$ and define $\hat{\bm{\mathrm{g}}}_{0}(\hat{y}):={\bm{\mathrm{g}}}(\hat{y})-{\bm{\mathrm{g}}}(\hat{y}_{0})$ for the corresponding pivot $\hat{y}_{0}\in\mathcal{Y}_{2}$. Then consider $q$ elements $y_{i}\in\mathcal{Y}_{1}$ and their correspondents $\hat{y}_{i}\in\mathcal{Y}_{2}$ such that the matrices:
${\bm{\mathrm{G}}}:=\begin{pmatrix}{\bm{\mathrm{g}}}_{0}(y_{1}),\ldots,{\bm{\mathrm{g}}}_{0}(y_{q})\end{pmatrix},\quad\hat{\bm{\mathrm{G}}}:=\begin{pmatrix}\hat{\bm{\mathrm{g}}}_{0}(\hat{y}_{1}),\ldots,\hat{\bm{\mathrm{g}}}_{0}(\hat{y}_{q})\end{pmatrix}$ (234)
have rank equal to $\mathrm{dim}(\Gamma_{1})$ and $\mathrm{dim}(\Gamma_{2})$, respectively. Then, we make use of these matrices with their transpose in 233, obtaining:
${\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ $=\beta\hat{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})$ (235)
$({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ $=\beta({\bm{\mathrm{G}}}^{\top})^{+}\hat{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})$ (Multiplying on the left for $({\bm{\mathrm{G}}}^{\top})^{+}$)
${\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ $=\beta{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})$ (236)
where in the last line we used $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}={\bm{\mathrm{P}}}_{\Gamma_{1}}$, since ${\bm{\mathrm{G}}}$ in 234 spans $\Gamma_{1}$, and we defined ${\bm{\mathrm{O}}}:=({\bm{\mathrm{G}}}^{\top})^{+}{\hat{\bm{\mathrm{G}}}}^{\top}$. Proceeding similarly but from $\hat{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})={\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ we obtain a similar equation to 236 :
${\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})=\beta^{-1}\hat{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})$ (237)
where we defined $\hat{\bm{\mathrm{O}}}:=(\hat{\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}$.
Let $k_{1}:=\mathrm{dim}(\Gamma_{1})$, $k_{2}:=\mathrm{dim}(\Gamma_{2})$, and $o:=\mathrm{rank}({\bm{\mathrm{O}}})$. Notice that, by definition, ${\bm{\mathrm{O}}}={\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{O}}}$, which means that:
$\mathrm{rank}({\bm{\mathrm{O}}})$ $=\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{O}}})$ (238)
$o$ $\leq\min(k_{1},o)$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$\implies o$ $\leq k_{1}$ (239)
Now, consider $\ell$ points, or sequences, ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that:
${\bm{\mathrm{F}}}_{1}:=\begin{pmatrix}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1}),\ldots,{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{1})\end{pmatrix},\quad{\bm{\mathrm{F}}}_{2}:=\begin{pmatrix}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{1}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2}),\ldots,{\bm{\mathrm{f}}}({\bm{\mathrm{x}}}_{\ell}\mathbin{\smallfrown}{\bm{\mathrm{q}}}_{2})\end{pmatrix},$ (240)
have rank equal to $\mathrm{dim}(\mathcal{F}_{1})$ and to $\mathrm{dim}(\mathcal{F}_{2})$, respectively. Therefore,by substituting ${\bm{\mathrm{F}}}_{1}$ and ${\bm{\mathrm{F}}}_{2}$ in 236 and in 237, we obtain:
${\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{F}}}_{1}$ $=\beta{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2}$ (241)
${\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2}$ $=\beta^{-1}\hat{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{F}}}_{1}\,.$ (242)
Then, by Lemma 17 it holds that:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{F}}}_{1})=\mathrm{dim}(\Gamma_{1})=k_{1},\quad\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2})=\mathrm{dim}(\Gamma_{2})=k_{2}\,.$ (243)
We use this in 241 to obtain the following:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{F}}}_{1})$ $=\mathrm{rank}(\beta{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2})$ (244)
$k_{1}$ $\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{O}}}),\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2})\big{)}$ (Using $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$)
$k_{1}$ $\leq\min(o,k_{2})$ (245)
$\implies$ $o\geq k_{1},\quad k_{2}\geq k_{1}\,.$ (246)
This shows that $k_{1}\leq o\leq k_{1}$, meaning that $\mathrm{rank}({\bm{\mathrm{O}}})=\mathrm{dim}(\Gamma_{1})$. Similarly, taking the rank of 242 we obtain a similar implication for $k_{1}$ and $k_{2}$ (as obtained for $k_{2}$ and $k_{1}$ for 241), that is:
$\mathrm{rank}({\bm{\mathrm{P}}}_{\Gamma_{2}}{\bm{\mathrm{F}}}_{2})$ $=\mathrm{rank}(\beta^{-1}\hat{\bm{\mathrm{O}}}{\bm{\mathrm{P}}}_{\Gamma_{1}}{\bm{\mathrm{F}}}_{1})$ (247)
$\implies$ $k_{2}\leq k_{1}\,,$ (248)
which shows that $k_{1}\leq k_{2}\leq k_{1}$, meaning that $\mathrm{dim}(\Gamma_{2})=\mathrm{dim}(\Gamma_{1})$. Moreover, the matrix ${\bm{\mathrm{O}}}$, being of rank $o=k_{1}=k_{2}$ defines an invertible transformation from $\Gamma_{2}$ to $\Gamma_{1}$. This concludes the proof. ∎
E.2 Tautologies
[…]
Definition 23 (Tautology).
We say that ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ is a tautology for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})\in\Theta$ if, for every ${\bm{\mathrm{s}}}\in\mathsf{Seq}(\mathcal{A})$ and all $y\in\mathcal{A}$, it holds:
$\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})=\log p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{q}}})\,.$ (249)
Next, we show that, for such tautologies, a “trivial” form of linear relational embedding holds.
Proposition 24 (Tautologies).
Let ${\bm{\mathrm{q}}}\in\mathsf{Seq}(\mathcal{A})$ be a tautology for the model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$, then ${\bm{\mathrm{f}}}$ linearly represents ${\bm{\mathrm{q}}}$ on $\mathcal{G}:=\mathrm{SIm}({\bm{\mathrm{g}}}_{0})$ with:
${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}\,.$ (250)
Proof.
From Definition 23, consider a pivot token $y_{0}\in\mathcal{A}$ and define ${\bm{\mathrm{g}}}_{0}:={\bm{\mathrm{g}}}(y)-{\bm{\mathrm{g}}}(y_{0})$ for every $y\in\mathcal{A}$. We have that:
$\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ $=\log\frac{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid{\bm{\mathrm{q}}})}{p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid{\bm{\mathrm{q}}})}$ (Take log of the ratio between $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y\mid\cdot)$ and $p_{{\bm{\mathrm{f}}},{\bm{\mathrm{g}}}}(y_{0}\mid\cdot)$)
$\log\frac{\exp{\bm{\mathrm{g}}}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}{\exp{\bm{\mathrm{g}}}(y_{0})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ $=\log\frac{\exp{\bm{\mathrm{g}}}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{q}}})}{\exp{\bm{\mathrm{g}}}(y_{0})^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{q}}})}$ (Write with the exponentional)
$\log{\exp{\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})}$ $=\log{\exp{\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{q}}})}$ (Use the definition of ${\bm{\mathrm{g}}}_{0}$)
${\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{q}}})$ (Remove log and exponential)
${\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{g}}}_{0}(y)^{\top}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}\,,$ (251)
where in the last line with denoted with ${\bm{\mathrm{a}}}_{\bm{\mathrm{q}}}:={\bm{\mathrm{f}}}({\bm{\mathrm{q}}})$. Consider $\ell$ tokens $y_{i}\in\mathcal{A}$ such that
${\bm{\mathrm{G}}}:=\begin{pmatrix}{\bm{\mathrm{g}}}_{0}(y_{1}),\ldots,{\bm{\mathrm{g}}}_{0}(y_{\ell})\end{pmatrix}$ (252)
spans $\mathcal{G}$. We use this and consider the following expression ofr the transpose:
${\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}$ (253)
$({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $=({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}$ (Multiply by pseudo-inverse of ${\bm{\mathrm{G}}}^{\top}$)
${\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{f}}}({\bm{\mathrm{s}}}\mathbin{\smallfrown}{\bm{\mathrm{q}}})$ $={\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{a}}}_{{\bm{\mathrm{q}}}}$ (254)
where we used the fact that $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathcal{G}}$. This shows the claim. ∎